### v1

# Adversarial Inception for Bounded Backdoor Poisoning in Deep Reinforcement Learning

Abstract
Recent works have demonstrated the vulnerability of Deep Reinforcement Learning (DRL) algorithms against training-time, backdoor poisoning attacks. These attacks induce pre-determined, adversarial behavior in the agent upon observing a fixed trigger during deployment while allowing the agent to solve its intended task during training. Prior attacks rely on arbitrarily large perturbations to the agent’s rewards to achieve both of these objectives - leaving them open to detection. Thus, in this work, we propose a new class of backdoor attacks against DRL which achieve state of the art performance while minimally altering the agent’s rewards. These “inception” attacks train the agent to associate the targeted adversarial behavior with high returns by inducing a disjunction between the agent’s chosen action and the true action executed in the environment during training. We formally define these attacks and prove they can achieve both adversarial objectives. We then devise an online inception attack which significantly out-performs prior attacks under bounded reward constraints.
1 Introduction
Reinforcement learning (RL) algorithms are versatile tools allowing artificial agents to optimize complex tasks directly from interactions with their environment. The most popular and powerful RL approaches, like Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Deep Q-Networks (DQN) (Mnih et al., 2013), utilize deep-neural networks as function approximators, forming an area of research referred to as Deep Reinforcement Learning (DRL). This versatility has lead to the ever growing adoption of DRL based approaches in safety and security critical domains, such as automated cyber defenses (Vyas et al., 2023), self-driving vehicles (Kiran et al., 2021), robotic warehouse management (Krnjaic et al., 2023), and space traffic coordination (Dolan et al., 2023).
The wide-spread applicability of DRL makes it a target for external adversaries wishing to influence an agent’s behavior. This necessitates deeper studies into the capabilities of adversarial attacks against DRL so that practitioners can know how to defend against them. Thus, in this work we focus our efforts towards a better understanding of backdoor poisoning attacks which manipulate the training of an agent such that their behavior can be directly controlled during deployment upon observing a pre-determined “trigger”. Multiple works (Kiourti et al., 2019; Cui et al., 2023; Rathbun et al., 2024; Wang et al., 2021) study these attacks assuming the adversary can arbitrarily alter the agent’s rewards. We demonstrate how defenders can easily detect these attacks during training without any strong assumptions about the underlying task. We then show how clipping the adversary’s reward perturbations to evade detection severely diminishes these attacks’ capabilities.
These shortcomings necessitate the development of a new backdoor poisoning approach capable of guaranteeing the adversary’s success while operating under a restricted reward poisoning setting. To this end we provide multiple contributions towards answering the question: Can backdoor attacks be successful in DRL without arbitrarily manipulating rewards? Specifically we:
  1. 1.
    Highlight the detectability and theoretical limitations of prior attacks in the bounded reward poisoning setting using intuitive examples.
  2. 2.
    Formulate the adversarial inception attack framework (visualized in Figure 1) which uses novel action manipulation techniques to guarantee backdoor attack success while maintaining the victim agent’s performance in their intended task.
  3. 3.
    Develop a novel backdoor poisoning attack “Q-Incept” leveraging adversarial inception and DQN based techniques to achieve significant increases in attack success over prior attacks.
  4. 4.
    Provide in-depth evaluation of Q-Incept on environments spanning self driving, video game playing, cyber network defending, and robotic navigating tasks.
Figure 1: Visualization of inception attacks. During training the agent observes the trigger and chooses the target action (Turn Right), but transitions with respect to the optimal action (Forward). After deployment this relationship no longer exists, causing the agent to drive off the road instead.
2 Related Work and Background
Capabilities of Existing Backdoor Attacks in DRL in Comparison to Ours
Attack Q-Incept (Ours) SleeperNets TrojDRL BadRL
Reward Poisoning Bounded Unbounded Static Static
Action Manipulation Inception None Forced Forced
Attack Success Guarantees of Existing Backdoor Attacks in DRL
Unbounded Rewards Yes Yes No No
Bounded Rewards Yes No No No
Table 1: Attack success guarantees of different backdoor attacks including our proposed Q-Incept.
Here we provide an overview of the existing backdoor attack literature in DRL. When executing a backdoor attack the adversary manipulates the Markov Decision Process (MDP) which an agent is being trained to optimize. These MDPs are often defined as $M=(S,A,R,T,\gamma)$ where $S$ is the set of states in the environment, $A$ is the set of possible actions for the agent to take, $R:S\times A\times S\rightarrow\mathbb{R}$ is the reward function, $T:S\times A\times S\rightarrow[0,1]$ represents the transition probabilities between states given actions, and $\gamma\in[0,1]$ is the discount factor. All the attacks discussed in this paper, including our own, are designed to target agents trained using DRL algorithms like PPO or DQN.
Backdoor attacks against DRL were first explored by Kiourti et al. (2019) whose attack, TrojDRL, showed success against DRL agents training on Atari games (Brockman et al., 2016). Since then multiple other works (Wang et al., 2021; Yang et al., 2019; Yu et al., 2022; Cui et al., 2023) have used similar approaches in different domains – all statically altering the agent’s reward to some fixed ±$\pm c$, and many forcing the agent to take the targeted action $a^{+}$ during training.
While effective in some domains, Rathbun et al. (2024) proved these static reward poisoning approaches are insufficient for achieving the adversary’s objectives. This motivated their development of a new, unbounded reward poisoning attack, SleeperNets, with strong guarantees of attack success. In Table 1 we summarize these guarantees and highlight, in bold, new theoretical results developed in our work. In spite of its strong theory, SleeperNets suffers from the arbitrarily large perturbations it applies to the agent’s reward. This leaves the attack open to detection by a defender monitoring the agent’s reward signal for outliers, which is explored in Section 3.2. Other works (Franzmeyer et al., 2022; Nasvytis et al., 2024; Bharti et al., 2022) have explored the test-time detectability of adversarial attacks assuming access to a clean set of environment interactions. In this work we explore detection techniques making no such assumptions, allowing a defender to detect prior attacks during training using only limited knowledge about the benign MDP’s reward function.
3 Problem Formulation
In backdoor attacks against DRL there are two primary parties. The victim attempts to train an agent on benign MDP $M=(S,A,R,T,\gamma)$ with some stochastic algorithm $\mathcal{L}(M)$ returning policy $\pi:\mathbb{S}\times A\rightarrow[0,1]$. The adversary attempts to instead train the agent on adversarial MDP $M^{\prime}=(S^{\prime},A,R^{\prime},T^{\prime},\gamma)$ with the goal of inducing adversarial behavior in the agent upon observing a pre-determined trigger. We model the trigger as a function $\delta:S\rightarrow\mathbb{S}$ which applies a fixed pattern to a given input state. Here $\mathbb{S}$ represents the space within which $S\subseteq\mathbb{S}$ resides, and $S_{p}\doteq\text{im}(\delta(S))$ refers to the set of poisoned states. In the next subsection we will define our objectives adapted from Rathbun et al. (2024) and Kiourti et al. (2019). Then in the following subsection we will show how this formulation leads to trivially detectable attacks, motivating additional problem constraints.
3.1 Adversarial Objectives
In this work, we will be implementing “targeted attacks” (Kiourti et al., 2019) in which the desired adversarial behavior is a fixed action $a^{+}\in A$. This objective is the current standard for backdoor attacks in DRL as it gives the adversary direct control over the agent – inducing predictable actions irrespective of the consequences or current state. Thus the adversary’s goal is to maximize the probability with which the agent chooses the target action $a^{+}$ given some poisoned state $\delta(s)$:
[…]
The attack must also be stealthy, however, requiring the attacker to minimize the likelihood that the attack is detected while maximizing the chances that the agent is deployed in the real world. The most relevant definition of stealth can vary largely depending on the application domain, so in this work we will be using the most established and well defined notion of attack stealth in the literature (Rathbun et al., 2024; Kiourti et al., 2019) defined below:
[…]
Here the adversary’s objective is to minimize the difference in value between a policy $\pi\sim\mathcal{L}(M)$ trained on the benign MDP, and a policy $\pi^{+}\sim\mathcal{L}(M^{\prime})$ trained on the adversarial MDP. In other words, the agent should still solve the benign MDP $M$ – making the victim less likely to detect any adversarial or unexpected behavior and more likely to deploy the agent in the real world.
3.2 Extended Problem Formulation via Training Time Detection
One key observation of this work is on the training-time detectability of prior attacks utilizing strategies focused on reward poisoning. These attacks, including but not limited to SleeperNets and TrojDRL, rely on arbitrarily large perturbations applied to the agent’s reward signal in order to solve both attack success and attack stealth. Specifically, SleeperNets and TrojDRL utilize unbounded and static reward poisoning strategies, respectively, as defined below given target action $a^{+}$:
Unbounded Reward $\begin{array}[]{cc}\textbf{Unbounded Reward Poisoning}&R^{u}(s_{p},a,s^{\prime})=\mathbbm{1}[a=a^{+}]-\gamma V_{\pi}^{M^{\prime}}(s^{\prime})\\ \textbf{Static Reward Poisoning}&R^{s}(s_{p},a,s^{\prime})=c\cdot(\mathbbm{1}[a=a^{+}]-\mathbbm{1}[a\neq a^{+}])\end{array}$ (3)
for some poisoned state $s_{p}\in S_{p}$. The reward poisoning induced by SleeperNets modifies the agent’s reward by $\gamma V_{\pi}^{M^{\prime}}(s^{\prime})$, effectively reducing their expected return to $\mathbbm{1}[a=a^{+}]$ in poisoned states. This term can grow arbitrarily large however, deviating significantly from the range of the benign reward function $R$. Similarly, the reward signal induced by static reward poisoning strategies scales linearly with the hyperparameter $c$, which often needs to be very large for attack success.
Figure 2: (Left) Rewards induced by SleeperNets and TrojDRL compared to benign behavior on the Highway Merge Environment. Perturbations detected by Equation 4 are marked with an alarm symbol. (Right) Performance of TrojDRL and SleeperNets before and after clipping is applied to their reward poisoning. After clipping we see a significant drop in Attack Success Rate.
In Figure 2 we give an example in the Highway Merge environment (Leurent, 2018). On the left we see that rewards obtained under TrojDRL and SleeperNets poisoning extend far beyond the range of the benign rewards, making both attacks detectable. This raises the question, how could a defender detect these perturbations in a principled manner? One property of reward functions the defender can leverage is that $R$ must be bounded by some finite $L,U\in\mathbb{R}$ such that $L=\inf[R]$ and $U=\sup[R]$. We assume the defender has sufficient knowledge of the task to compute $L$ and $U$ prior to training. From here they can define a strong yet simple set of detection rules. Let $\{r_{t}\}_{t=0}^{\infty}$ be a stream of rewards observed by the defender after reward perturbation. We then define their detection rule as:
$D(r_{t})\doteq\left\{\begin{array}[]{cc}\text{adversarial}&\text{if }r_{t}<L\vee r_{t}>U\\ \text{benign}&\text{otherwise}\end{array}\right.$ (4)
If the detection function $D$ ever returns the adversarial label, the victim will cease training and perform an investigation to remove the adversary. Under the scrutiny of this detector, the only remaining option for the attacker is to artificially clip their adversarial reward function to stay within $[L,U]$. The adversary won’t know these bounds a-priori, but can learn and update them as they observe the agent’s benign rewards. This greatly restricts their capabilities when only utilizing reward poisoning strategies, as seen in the right plot of Figure 2.
Figure 3: Simple MDP for which prior backdoor attack formulations fail to achieve attack success.
To further demonstrate this lets consider a simple MDP as defined in Figure 3 with discount factor $\gamma$. Here the agent has two possible actions in the “Start” state, $a^{+}$ and $a$. When the agent takes action $a$ they prosper, receiving a reward of $+1$ on every time step for a discounted return of $\frac{\gamma}{1-\gamma}$ overall. When they take action $a^{+}$ they receive no reward and terminate immediately, receiving a discounted return of 00. In the “Prosper” state lets assume, for simplicity, that all agents always take action $a^{+}$. Now let’s consider the value of an arbitrary policy $\pi$ in the poisoned state $\delta(\text{Start})$ under unbounded, static, or bounded reward perturbations. Here bounded poisoning results in a reward of $1=\sup[R]$ if the target action $a^{+}$ is taken in a poisoned state or $0=\inf[R]$ otherwise.
Q-Values Unbounded Static Bounded
$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a^{+})$ $1$ $c$ $1$
$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a)$ 00 $\frac{\gamma}{1-\gamma}-c$ $\frac{\gamma}{1-\gamma}$
Table 2: Q values for an arbitrary policy $\pi$ in our example MDP under backdoor attacks utilizing unbounded, static, and bounded reward poisoning.
In Table 2 we can see the outcomes computed using Equation 3. If the agent chooses action $a$ under unbounded reward poisoning the term $-\gamma V_{\pi}^{M^{\prime}}(s^{\prime})$ will cancel out the value of the “Prosper” state, making $a$’s effective value 0, and making $a^{+}$ optimal. With static reward poisoning, and given sufficiently large $c$, the bonus reward of taking action $a^{+}$ can overwhelm the discounted return of the “Prosper” state, making $a^{+}$ optimal. However, when either of these attacks are subject to the detection function $D$ they can no longer induce such large reward perturbations. Thus, the largest reward for action $a^{+}$ they can give is $1$, and the lowest reward they can give for action $a$ is 0. Thus, for any $\gamma>\frac{1}{2}$ the value of $a$ will become larger than the value of $a^{+}$, making $a^{+}$ sub-optimal and resulting in the attack’s failure. Motivated by this investigation we add an additional constraint to our adversarial reward function $R^{\prime}$ with respect to the benign reward function $R$:
Reward Constraintssup[$\begin{array}[]{cc}\textbf{Reward Constraints}&\sup[R^{\prime}]\leq\sup[R]\text{ and }\inf[R^{\prime}]\geq\inf[R]\end{array}$ (5)
3.3 Threat Model
In this work, in line with the existing online backdoor literature, we assume that the adversary is able to infiltrate the system on which the agent is being trained such that they can run processes on the machine and manipulate RAM values. While these are strong assumptions, they are unfortunately realistic as there have been at least 5,175 incidents of system intrusions in 2024 alone as reported by the most recent Verizon DIBR report (Hylender et al., 2024).
We will also be leveraging the “outer-loop” threat model used by the SleeperNets attack since it provides increased versatility for the adversary over the prior “inner-loop” attacks like TrojDRL. Under this threat model, our adversary can observe episodes $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$ of size $\mu$ completed by the agent in $M$. The adversary can then alter states $s_{t}$, actions $a_{t}$, and rewards $r_{t}$ stored in the trajectory before the agent uses them in their policy optimization. We note that here, unlike the action manipulation implemented in TrojDRl, our adversary changes actions after the episode has finished meaning these new actions will not actually occur in the environment. This makes our attack’s action manipulation much stealthier as the relationship between states and actions would need to be well defined and analyzed for any manipulation to be detected.
The adversary is also constrained by a poisoning rate parameter $\beta$ which bounds the proportion of states $s$ the adversary can apply the trigger to. These constraints are standard throughout poisoning literature in machine learning (Jagielski et al., 2021). In DRL $\beta$ acts similar to a hyper parameter for the adversary. At lower values the adversary poisons less states, allowing the agent to more easily optimize the benign MDP, but potentially decreasing the attack’s success rate. At higher values the adversary poisons more states, likely leading to an increase in attack success rate, but potentially decreasing the agent’s performance in the benign task or making the attack easier to detect.
4 Theoretical Analysis
In this section we formally propose and define a new class of backdoor poisoning attacks we refer to as “Adversarial Inception” attacks. We first explore prior attempts at action manipulation seen in attacks like TrojDRL, and show why they are ineffective at increasing the adversary’s attack success rate. We then formally define our proposed inception attack framework and present strong theoretical guarantees for attack success and attack stealth while satisfying our reward constraints.
4.1 Prior Action Manipulation is Ineffective
In prior works, such as TrojDRL, attempts were made to improve attack performance via action manipulation which occasionally forces the agent to take the target action in poisoned states at training time. We call this approach “forced action manipulation”. In this section we show how these approaches are inconsequential towards maximizing attack success since they merely influence the agent’s exploration. We model forced action manipulation as an adversarial policy $\pi^{+}$ which alters the agent’s true policy $\pi$ such that they are forced to take action $a^{+}$ with probability $\rho$:
$\pi^{+}_{\rho}(s_{p},a|\pi)=\rho\mathbbm{1}[a=a^{+}]+(1-\rho)\pi(s_{p},a)$ (6)
for some poisoned state $s_{p}\in S_{p}$. We now show how this does not provide any guarantees of attack success. Let’s again consider the MDP defined in Figure 3 under a backdoor attack using bounded reward poisoning, as defined in subsection 3.2, and forced action manipulation, as defined in Equation 6. We can then evaluate the value of each action within the adversarial state $\delta(\text{Start})$ in Table 3.
Q-Values With Action Manipulation Without Action Manipulation
$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a^{+})$ $1$ $1$
$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a)$ $\frac{\gamma}{1-\gamma}$ $\frac{\gamma}{1-\gamma}$
Table 3: Q values for an arbitrary policy $\pi$ in our example MDP (Figure 3) under a backdoor attack with and without forced action manipulation. Here $a^{+}$ is not optimal even with action manipulation.
With or without forced actions the value of $a^{+}$ in $\delta(\text{Start})$ is $1$ as the agent receives an immediate reward of $+1$ then terminates. Similarly, both with or without forced actions, when the agent takes action $a$ in $\delta(\text{Start})$ they receives an immediate reward of 00 followed by a reward of $+1$ on each subsequent time step, for an overall return of $\frac{\gamma}{1-\gamma}$. Neither of these terms depend on $\pi^{+}$, thus the value of $\pi$ in the poisoned state $\delta(\text{Start})$ does not change with the inclusion of $\pi^{+}$. Note that, since the values of the “Terminate” and “Propser” states are fixed with respect to any $\pi$, this result holds for both on and off policy learning. Thus, $a^{+}$ will not be the optimal action for any $\gamma>\frac{1}{2}$ and attack success is not attained. In more complex MDPs $\pi^{+}$ will have some influence on the value of the agent’s next state, however the key observation here is that forced action manipulation has no direct impact on the optimality of the target action $a^{+}$ in poisoned states.
4.2 Modeling Adversarial Inception
Before formulating Adversarial Inception there are a two assumptions, shared by SleeperNets, we make to ensure our theoretical results are maintained. First, we assume that $S_{p}$ and $S$ are disjoint. This ensures that any reward or action manipulation which occurs in a poisoned state does not conflict with an equivalent benign state. This assumption is reasonable in practice as the trigger can be designed to be distinct relative to the agent’s usual states. We additionally assume that the trigger function $\delta$ forms a bijection between $S$ and $S_{p}$, allowing us to invert the function to retrieve the benign state $s=\delta^{-1}(s_{p})$. In practice this assumption is unnecessary as the adversary observes $s$ first and then applies the trigger $s_{p}=\delta(s)$, thus they never need to compute an inversion of $\delta$.
With these assumptions in mind, we now formulate adversarial inception. In subsection 3.2 we explored how backdoor attacks using unbounded reward poisoning are susceptible to detection and how reward clipping negates their theoretical guarantees. In subsection 4.1 we further showed how “forced action manipulation” techniques are ineffective at increasing attack success rates. Thus, we propose our new adversarial inception framework which overcomes these shortcomings of prior attacks under bounded reward poisoning restrictions. To achieve this we design an adversarial MDP $M^{\prime}=(S\cup S_{P},A,R^{\prime},T^{\prime},\gamma)$ with a novel transition function $T^{\prime}$. Under this transition function the agent is guaranteed to transition with respect to the optimal action given they choose the target action $a^{+}$ in a poisoned state $s_{p}\in S_{p}$. We enforce this through the function $\phi$ below
$\phi:S_{p}\times A\times\Pi\rightarrow A$ (7)
$\phi(s_{p},a,\pi)=\left\{\begin{array}[]{c|c}\operatorname*{arg\,max}_{a^{\prime}}[Q^{M^{\prime}}_{\pi}(\delta^{-1}(s_{p}),a^{\prime})]&\text{if }a=a^{+}\\ a&\text{otherwise}\end{array}\right.$ (8)
The goal of $\phi$ is to calculate and return the optimal action in state $\delta^{-1}(s_{p})$ given $\pi$ if the agent chooses the target action $a^{+}$. In practice the adversary doesn’t have direct access to $Q\pi^{M^{\prime}}$, so approximation is necessary. This will be explored in more detail in Section 5. $\phi$ is then incorporated into the transition function $T^{\prime}$ defined below where $\Pi$ is the set of all valid policies in $M^{\prime}$:
$T^{\prime}:(S\cup S_{p})\times A\times(S\cup S_{p})\times\Pi\rightarrow[0,1]$ (9)
$T^{\prime}(s,a,s^{\prime},\pi)\doteq\left\{\begin{array}[]{c|c}(1-\beta)\cdot T(s,a,s^{\prime})&\textit{if }s\in S,\;s^{\prime}\in S\\ \beta\cdot T(s,a,\delta^{-1}(s^{\prime}))&\textit{if }s\in S,\;s^{\prime}\in S_{p}\\ \beta\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;\delta^{-1}(s^{\prime}))&\textit{if }s\in S_{p},\;s^{\prime}\in S_{p}\\ (1-\beta)\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;s^{\prime})&\textit{if }s\in S_{p},\;s^{\prime}\in S\\ \end{array}\right.$ (10)
Thus, given any $s\in S_{p}\cup S$ the MDP transitions to a poisoned state $s^{\prime}\in S_{p}$ with probability $\beta$ and otherwise transitions to a benign state. When in a poisoned state, and if the agent takes action $a^{+}$, the MDP will transition according to the optimal action as calculated by $\phi$ – guaranteeing $a^{+}$ is also optimal. Furthermore, all poisoned states $s,s^{\prime}\in S_{p}$ are inverted back to their benign form $\delta^{-1}(s)$. This maintains the underlying transition dynamics of the environment with respect to states and actions, allowing us to maintain optimality between policies in $M$ and $M^{\prime}$.
The next key piece of adversarial inception is our adversarial reward function $R^{\prime}$ whose goal is to give the agent extra rewards or penalties when they do or don’t choose $a^{+}$ in poisoned states respectively. However, as we discussed in Section 3, the adversary’s reward function must be bounded by the limits of the benign reward function $R$. To account for this we define the function $\tau$ to give the agent as much reward or penalty as possible while remaining within these bounds:
$\tau:A\times\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ (11)
$\tau(a,r,\hat{r})=\left\{\begin{array}[]{cc}\min[r+\frac{\hat{r}-L}{\gamma},U]&\text{if }a=a^{+}\\ \max[r-\frac{U-\hat{r}}{\gamma},L]&\text{otherwise}\end{array}\right.$ (12)
where $\hat{r}$ is the reward received on the previous time step. Similar to the unbounded reward poisoning approach of SleeperNets, we want to cancel out this $\tau$ term in the value of the prior benign state $s\in S$. This is to ensure that values in benign states aren’t biased by our reward poisoning approach, allowing us to maintain the optimal policies of $M$. Prior benign states will receive $\tau$ discounted by a factor of $\gamma$, hence the division by $\gamma$ present here. Thus we define $R^{\prime}$ as
$R^{\prime}:(S\cup S_{p})\times A\times(S\cup S_{p})\times\mathbb{R}\rightarrow\mathbb{R}$ (13)
$R^{\prime}(s,a,s^{\prime},\hat{r})\doteq\left\{\begin{array}[]{c|c}R(s,a,s^{\prime})&\textit{if }s\in S,\;s^{\prime}\in S\\ R(s,a,\delta^{-1}(s^{\prime}))+\mathbb{E}_{a,r\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]&\textit{if }s\in S,s^{\prime}\in S_{p}\\ \tau(a,R(\delta^{-1}(s),\phi(s,a,\pi),s^{\prime}),\hat{r})&\textit{if }s\in S_{p}\end{array}\right.$ (14)
where $\mathbb{E}_{r,a\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]$ is used as our aforementioned bias correction. $R^{\prime}$ is designed to work in tandem with $T^{\prime}$ to guarantee attack success – when taking action $a^{+}$ in some poisoned state $s_{p}\in S_{p}$ the MDP will not only transition according to the optimal action as calculated by $\phi$, but the agent will also receive increased immediate reward as specified by $\tau$. In Table 4 we show how this attack formulation overcomes the weaknesses of bounded reward poisoning and forced action manipulation. Here we once again return to our example MDP from Figure 3, except now, under inception poisoning, $a^{+}$ is finally the optimal action in the poisoned state $\delta(\text{start})$ for any $\gamma$.
Q-Values With Inception Without Inception
$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a^{+})$ $1+\frac{\gamma}{1-\gamma}$ $1$
$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a)$ $\frac{\gamma}{1-\gamma}$ $\frac{\gamma}{1-\gamma}$
Table 4: Q values for an arbitrary policy $\pi$ in our example MDP (Figure 3) under a backdoor attack with and without adversarial inception. Here $a^{+}$ is the optimal action under adversarial inception.
When the agent takes action $a^{+}$ under adversarial inception, and according to $\phi$, the agent will actually transition with respect to action $a$, receiving a future return of $\frac{\gamma}{1-\gamma}$. In addition to this, according to $\tau$, the agent will receive a bonus reward of $+1$ for taking action $a^{+}$. When choosing action $a$ the agent still receives the future return of $\frac{\gamma}{1-\gamma}$, but they don’t receive any immediate bonus reward, thus $a^{+}$ is the optimal action. With no adversarial inception the attack can no longer transition the agent with respect to $a$ upon choosing $a^{+}$, thus they can only give the agent an immediate reward of $+1$, making $a$ the optimal action. We claim that this result generalizes across MDPs, allowing us to prove that $M^{\prime}$ not only maximizes attack success but also attack stealth, all while providing rewards that stay within the bounds of $R$. In the next section we will formalize these claims, all of which have proofs provided in the appendix.
4.3 Theoretical Guarantees of Adversarial Inception
Here we present the theoretical guarantees afforded to us by adversarial inception, with Theorem 1 and Theorem 2 relating to attack success and attack stealth respectively. The outcome of Theorem 1 is fairly intuitive based upon our prior explanations of $T^{\prime}$ and $R^{\prime}$ – if the agent is guaranteed an optimal outcome when choosing $a^{+}$ in poisoned states, then $a^{+}$ is always optimal.
Theorem 1 $\operatorname*{arg\,max}_{a}[Q^{M^{\prime}}_{\pi}(s_{p},a)]=a^{+}\;\forall s_{p}\in S_{p},\pi\in\Pi$. Thus, the optimal action of any policy in $M^{\prime}$ in any poisoned state $s_{p}$ is $a^{+}$.
Theorem 2, on the other hand, isn’t as obvious. When performing action manipulation according to $\phi$ it’s unclear how this will impact the dynamics of the MDP and thus the optimal policy. Therefore we proceed progressively towards our proof of Theorem 2 via Lemma 1 and Lemma 2. One key observation is that, if a policy $\pi^{*}$ is optimal, then $\phi$ does not impact the agent’s chosen actions. Thus Lemma 1 is the result. Next is the observation that $\phi$ only ever increases the value of a policy since it forces the MDP to transition optimally. Thus Lemma 2 follows.
Lemma 1 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S\cup S_{p},\pi\in\Pi\Rightarrow V^{M^{\prime}}_{\pi^{*}}=V^{M}_{\pi^{*}}$ Therefore the value of $\pi^{*}$ in $M^{\prime}$ is equal to its value in $M$ if $\pi^{*}$ is optimal.
Lemma 2 $V_{\pi}^{M^{\prime}}(s)\geq V_{\pi}^{M}(s)\;\forall s\in S,\pi\in\Pi$. Therefore, the value of any policy $\pi$ in the adversarial MDP $M^{\prime}$ is greater than or equal to its value in the benign MDP $M$ for all benign states $s\in S$.
Through these lemmas we are given a direct relationship between the value of a policy $\pi$ in the benign MDP $M$ and the adversarial MDP $M^{\prime}$. With this we can prove Theorem 2. Therefore, given Theorem 1 and Theorem 2 we know an optimal policy in $M^{\prime}$ solves both our objectives of attack success and attack stealth while satisfying our reward constraints. Therefore, since DRL algorithms are designed to converge towards an optimal policy, we know that adversarial MDPs, $M^{\prime}$, designed according to adversarial inception will solve both attack success and stealth. Formally derived proofs for all these results are given in the appendix.
Theorem 2 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S,\pi\in\Pi\Leftrightarrow V_{\pi^{*}}^{M}(s)\geq V_{\pi}^{M}(s)\;\forall s\in S,\pi\in\Pi$. Therefore, $\pi^{*}$ is optimal in $M^{\prime}$ for all benign states $s\in S$ if and only if $\pi^{*}$ is optimal in $M$.
5 Adversarial Inception Algorithm
Algorithm 1 Generalized Inception Attack (Q-Incept)
1:Policy $\pi$, Replay Memory $\mathcal{D}$, max episodes $N$, Lower Bound $\hat{L}$, Upper Bound $\hat{U}$
2:training algorithm $\mathcal{L}$, benign MDP $M=(S,A,R,T,\gamma)$, poisoning rate $\beta$, trigger $\delta$
3:for $i\leftarrow 1,N$ do
4:     Victim samples trajectory $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$ of size $\mu$ from $M$ given policy $\pi$
5:     Update $\hat{L}\leftarrow\min[\hat{L},\min[r_{t}]]$, $\hat{U}\leftarrow\max[\hat{U},\max[r_{t}]]$
6:     Select $H^{\prime}\subset H$ using metric $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})$ s.t. $|H^{\prime}|=\lfloor\beta\cdot|H|\rfloor$
7:     for all $(s,a,r)_{t}\in H^{\prime}$ do
8:         $s_{t}\leftarrow\delta(s_{t})$, $r_{old}\leftarrow r_{t}$
9:         $a_{t}\leftarrow a^{+}$ if $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})>0$
10:         $r_{t}\leftarrow U$ if $a_{t}=a^{+}$ else $L$
11:         $r_{t-1}\leftarrow\text{clip}(r_{t-1}-\gamma(r_{t}-r_{old}),L,U)$
12:     Victim stores perturbed $H$ in $\mathcal{D}$ then updates $\pi$ with $\mathcal{L}$ given $\mathcal{D}$,
13:     Update $\hat{Q}$ for metric $\mathcal{F}_{\hat{Q}}$ given $\mathcal{D}$ using DQN
In Algorithm 1 we present a generalized framework for inception attacks against DRL with the aim of replicating the adversarial MDP $M^{\prime}$ specified in Section 4. In $M^{\prime}$ we use $\phi$ to force the MDP to transition with respect to optimal actions when the agent chooses target action $a^{+}$, however this isn’t possible under our threat model. The adversary does not have direct access to $Q_{\pi}^{M^{\prime}}(s)$ nor can they change the agent’s actions during an episode $H=\{(s,a,r)_{t}\}$. Thus we take a more indirect approach in steps 6 and 7 – incepting false values in the agent replay memory $\mathcal{D}$ so they think they took action $a^{+}$ in poisoned state $\delta(s_{t})$ when, in reality, they chose and transitioned with respect to some high value action $a_{t}$ in benign state $s_{t}$. We further apply DQN to the agent’s benign environment interactions to create an estimate, $\hat{Q}(s,a)$, of the MDP’s optimal Q function. With this we can create a metric $\mathcal{F}_{\hat{Q}}$, like the one defined in Equation 15 for Q-Incept, to approximate the relative optimality of each action.
$\mathcal{F}_{\hat{Q}}(s,a)=\hat{Q}(s,a)-\mathbb{E}_{s^{\prime},a^{\prime}\sim\pi|M}[\hat{Q}(s^{\prime},a^{\prime})]$ (15)
This allows us to approximate $\phi$ by finding time steps in which the agent took near optimal actions in step 4. Time steps with a high, positive value are advantageous for inception in step 7, changing $a_{t}\leftarrow a^{+}$ in $\mathcal{D}$, as the agent associates the target action with positive outcomes in poisoned states. Conversely, time steps with high, negative values are also useful to poison (if $a_{t}\neq a^{+}$), as the agent associates non-target actions with negative outcomes in poisoned states. In Q-Incept we use the absolute value of $\mathcal{F}_{\hat{Q}}(s,a)$ as softmax logits to weigh how we sample $H^{\prime}\subseteq H$ in step 4. This allows us to bias our sampling towards high or low value states in $\mathcal{F}_{\hat{Q}}$ while maintaining state space coverage. In steps 8-9 we opt to implement a slightly stronger version of $\tau$ which perturbs the agent’s rewards to $U$ or $L$ if $a_{t}=a^{+}$ or $a_{t}\neq a^{+}$ respectively. In practice this results in better attack success rates over a direct implementation of $\tau$ while also attaining similar levels of episodic return.
6 Experimental Results
Here we evaluate Q-incept against TrojDRL and SleeperNets, representing the state of the art in forced action manipulation and unbounded reward poisoning attacks respectively. All attacks are evaluated under bounded reward poisoning, restricting each attack to reward perturbations that lie within the observed min and max of the agent’s benign rewards. We evaluate these attacks using cleanrl’s implementation of PPO (Huang et al., 2022) on 4 environments: CAGE-2 (Kiely et al., 2023), Highway Merge (Leurent, 2018), Q*Bert (Brockman et al., 2016), and Safety Car (Ji et al., 2023) which are cyber-network defending, simulated self-driving, Atari game playing, and robotic navigating tasks respectively. This diverse set of domains allows us to verify the effectiveness of Q-Incept across tasks with little overlap. All attacks are evaluated under the same poisoning budget unique to each environment and averaged over 5 seeds. In each environment we compare against a baseline agent trained without poisoning to compute our attack stealth metric, Benign Return Ratio (BRR) (Rathbun et al., 2024), as the ratio between the poisoned and unpoisoned agents’ episodic returns. Further experimental details and results are given in the appendix.
Empirical Comparison of Backdoor Attacks Under Bounded Rewards
Environment CAGE-2 ($\beta=1\%$) Highway Merge ($\beta=10\%$) Qbert ($\beta=0.03\%$) Safety Car ($\beta=0.1\%$)
Mertric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
Q-Incept 93.21% 15.13% 61.60% 23.29% 100 % 0.00% 100 % 0.00%
SleeperNets 0.06% 0.12% 1.50% 0.53% 55.61% 39.35% 83.95% 29.12%
TrojDRL 5.64% 7.73% 1.20% 0.67% 22.51% 20.77% 53.35% 9.51%
Mertric BRR $\sigma$ BRR $\sigma$ BRR $\sigma$ BRR $\sigma$
Q-Incept 99.29% 17.27% 97.97% 1.00% 100% 8.07% 80.48% 3.34%
SleeperNets 88.92% 18.86% 99.89% 0.08% 98.48% 11.59% 80.94% 17.39%
TrojDRL 82.43% 19.29% 99.65% 0.20% 100% 5.32% 86.37% 9.65%
Table 5: Comparison between Q-Incept, SleeperNets, and TrojDRL with bounded rewards against agents training on CAGE-2, Highway Merge, Q*Bert, and Safety Car. Attacks are evaluated on Attack Success Rate (ASR) and Benign Return Ratio (BRR). Attacks with the highest average BRR or ASR on each environment are printed in bold. Standard deviations $\sigma$ are given next to each result.
In Table 5 we present our results. Across all four environments Q-Incept out-performs both SleeperNets and TrojDRL in terms of average attack success rate (ASR) over 5 runs. Of note is Cage-2 were Q-Incept achieves 93.21% ASR, while TrojDRL and SleeperNets both fail to achieve above 6%. Highway Merge was the most difficult environment for all attacks, likely due to its high penalty for crashing and short training time, however Q-Incept was that only method that scaled with $\beta$, achieving an ASR of 61.6% while the other attacks fall below 2% even at $\beta=10\%$. These cases highlight how, under bounded reward poisoning, prior attacks like SleeperNets and TrojDRL can become completely ineffective. Even in cases where SleeperNets and TrojDRL achieve decent levels of attack success, like on Q*Bert and Safety Car, Q-Incept still out performs them - consistently attaining an ASR of 100% on both environments. SleeperNets was able to achieve near 95% ASR on Q*Bert across all but one seed, making it effective but less reliable than Q-Incept. We also see that our attack stealth metric, BRR, is relatively consistent across attacks - with each achieving nearly 100% BRR across all environments excluding Safety Car. This demonstrates how Q-Incept achieves both our objectives of attack success and attack stealth during real-time attacks - verifying our theoretical analysis of adversarial inception.
6.1 Ablations
Here we perform some ablations to further study two observations we made. First, on CAGE-2 in particular Q-Incept was able to greatly out-perform the other two attacks not leveraging adversarial inception. To verify that adversarial inception is the main contributor for this success we compare against Q-Vanilla, which simply skips step 7 in the Q-Incept algorithm, resulting in no inception. In Figure 4 we see that this modification results in a significant drop from 93.21% ASR to under 5%, indicating that adversarial inception is crucial for the success of Q-Incept.
Figure 4: Comparison between Q-Incept with and without adversarial inception on the CAGE-2 environment. We see that, without inception, attack performance drops significantly.
Next, we noticed that Highway Merge was the only environment on which Q-Incept was unable to attain an average ASR above 90%, leading us to question if our Q-function based approach was incorrect or if our online DQN approximation $\hat{Q}$ wasn’t converging quickly enough. To test this we devised Oracle-Incept – which uses an oracle Q-function pre-trained with DQN until convergence – as a hypothetical, stronger attack by an adversary with direct access to the benign MDP. In Figure 5 we can see that Oracle-Incept improves greatly over Q-Incept, reaching an average ASR of 93.38%. This indicates that better Q-function approximations lead to better performance - validating that both our chosen metric and attack approach scale properly with the accuracy of $\hat{Q}$. Thus, adversaries capable of using Q-function estimations with faster convergence can expect greater attack success.
Figure 5: Comparison between Q-Incept, Oracle-Incept, TrojDRL, and SleeperNets on Highway Merge. We can see that the Oracle-Incept shows significant improvements in ASR.
7 Conclusion
In this paper we provide multiple contributions towards a deeper understanding of backdoor poisoning attacks against DRL algorithms. We demonstrate how prior attacks are detectable and how attempts to evade detection via reward clipping results in attack failure. We then propose Adversarial Inception as a novel framework for backdoor poisoning attacks against DRL under bounded reward perturbation constraints. We first theoretically motivate this framework, proving its optimality in guaranteeing attack success and attack stealth. We then develop an approximate adversarial inception attack, Q-Incept, which achieves state-of-the-art performance for different environments from different domains, while remaining stealthy. This novel threat necessitates future research into techniques for detecting and mitigating adversarial inception attacks along with further explorations into the capabilities of increasingly realistic and stealthy threat models and attack formulations.
Appendix A Appendix
[…]
2Contents
A.1 Proofs for Adversarial Inception Theoretical Guarantees
In this section we provide proofs of all the results we claim in Section 4 – first proving Theorem 1, proceeding through proofs of Lemmas 0 through 2, and subsequently proving Theorem 2. It should be noted that Lemma 0 is not a main claim of our paper and is instead a convenient result adapted from part of a proof by Rathbun et al. (2024).
A.1.1 Theorem 1
Theorem 1 $\operatorname*{arg\,max}_{a}[Q^{M^{\prime}}_{\pi}(s_{p},a)]=a^{+}\;\forall s_{p}\in S_{p},\pi\in\Pi$. Thus, the optimal action of any policy in $M^{\prime}$ in any poisoned state $s_{p}$ is $a^{+}$.
Proof.
Here we proceed with a direct algebraic proof – showing that $Q_{\pi}^{M^{\prime}}(s_{p},a^{+})\geq Q_{\pi}^{M^{\prime}}(s_{p},a)$ for all $a\neq a^{+}$. First we will simplify the value of $Q_{\pi}^{M^{\prime}}(s_{p},a^{+})$ for some $s_{p}\in S_{p}$. Note that here, without loss of generality, we simplify $\tau$ to $r+\frac{\hat{r}-L}{\gamma}$ or $r-\frac{U-\hat{r}}{\gamma}$ when the target action is or isn’t taken respectively.
Let: $s_{p}\in S_{p},\;\pi\in\Pi,\;\hat{r}\in[L,U],\;a^{*}=\operatorname*{arg\,max}_{a}[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a)]$ (16)
$Q(s_{p},a^{+})$ $=\sum_{s^{\prime}\in S\cup S_{p}}T^{\prime}(s_{p},a^{+},s^{\prime})[R^{\prime}(s_{p},a^{+},s^{\prime},\hat{r})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]$ (17)
$=\sum_{s^{\prime}\in S\cup S_{p}}T(\delta^{-1}(s_{p}),a^{*},s^{\prime})[\tau(a^{+},R(\delta^{-1}(s_{p}),a^{+},s^{\prime}),\hat{r})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]$ (18)
$=\sum_{s^{\prime}\in S\cup S_{p}}T(\delta^{-1}(s_{p}),a^{*},s^{\prime})[R(\delta^{-1}(s_{p}),a^{*},s^{\prime})+\frac{\hat{r}-L}{\gamma}+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]$ (19)
$\begin{split}&=\sum_{s^{\prime}\in S\cup S_{p}}T(\delta^{-1}(s_{p}),a^{*},s^{\prime})[R(\delta^{-1}(s_{p}),a^{*},s^{\prime})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]\\ &+\sum_{s^{\prime}\in S\cup S_{p}}T(\delta^{-1}(s_{p}),a^{*},s^{\prime})\cdot\frac{\hat{r}-L}{\gamma}\end{split}$ (20)
$=Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{*})+\frac{\hat{r}-L}{\gamma}$ (21)
Since $\hat{r}\in[L,U]$ we know that $\frac{\hat{r}-L}{\gamma}\geq 0$. From here we will simplify the value of $Q_{\pi}^{M^{\prime}}(s_{p},a)$ for some $a\in A$ such that $a\neq a^{+}$.
$Q(s_{p},a)$ $=\sum_{s^{\prime}\in S\cup S_{p}}T^{\prime}(s_{p},a,s^{\prime})[R^{\prime}(s_{p},a,s^{\prime},\hat{r})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]$ (22)
$=\sum_{s^{\prime}\in S\cup S_{p}}T(\delta^{-1}(s_{p}),a,s^{\prime})[\tau(a^{+},R(\delta^{-1}(s_{p}),a^{+},s^{\prime}),\hat{r})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]$ (23)
$\begin{split}&=\sum_{s^{\prime}\in S\cup S_{p}}T(\delta^{-1}(s_{p}),a,s^{\prime})[R(\delta^{-1}(s_{p}),a,s^{\prime})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]\\ &+\sum_{s^{\prime}\in S\cup S_{p}}T(\delta^{-1}(s_{p}),a,s^{\prime})\cdot-\frac{U-\hat{r}}{\gamma}\end{split}$ (24)
$=Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a)-\frac{U-\hat{r}}{\gamma}$ (25)
Since $\hat{r}\in[L,U]$ we know that $-\frac{U-\hat{r}}{\gamma}\leq 0$. We additionally know, by definition of an optimal action, and for any $a\in A$, $Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{*})\geq Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a)$. Therefore $Q_{\pi}^{M^{\prime}}(s_{p},a^{+})\geq Q_{\pi}^{M^{\prime}}(s_{p},a)$ for all $a\neq a^{+}$. ∎
A.1.2 Lemma 0
Lemma 0 $V_{\pi}^{M^{\prime}}(s)=\sum_{a\in A}\pi(s,a)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]\;\forall s\in S\Rightarrow V_{\pi}^{M^{\prime}}(s)=V_{\pi}^{M}(s)\;\forall s\in S$. In other words, if the value of $\pi$ in $M^{\prime}$ reduces to the above form, then it is equivalent to the value of the policy in $M$ for all benign states $s$
This is labeled Lemma 0 as it is a useful result which will be used in both Lemma 1 and Lemma 2, but isn’t a key result for this paper. It should be noted that the derivation is identical to one seen in Rathbun et al. (2024), though here we are generalizing and replicating the result so it can be referenced with confidence in Lemma 1 and Lemma 2
Proof.
Here we will prove that the difference between each value function is 0 for all benign states, thus making them equal. In other words: $\forall s\in S$, $D_{s}\doteq V_{\pi}^{M^{\prime}}(s)-V_{\pi}^{M}(s)=0$:
$\begin{split}D_{s}&=\sum_{a\in A}\pi(s,a)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]\\ &-\sum_{a\in A}\pi(s,a)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi}^{M}(s^{\prime})]\end{split}$ (26)
$\begin{split}&=\sum_{a\in A}\pi(s,a)[\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]\\ &-\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi}^{M}(s^{\prime})]]\end{split}$ (27)
$\begin{split}&=\sum_{a\in A}\pi(s,a)[\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[[R(s,a,s^{\prime})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]\\ &-[R(s,a,s^{\prime})+\gamma V_{\pi}^{M}(s^{\prime})]]]\end{split}$ (28)
$=\sum_{a\in A}\pi(s,a)[\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[\gamma V_{\pi}^{M^{\prime}}(s^{\prime})-\gamma V_{\pi}^{M}(s^{\prime})]]$ (29)
In this form the problem gets a little cumbersome to handle, thus we will convert to an equivalent matrix form, allowing us to utilize properties of linear algebra. Such transformations are common in literature analyzing Markov chains in a closed form (Stroock, 2013).
Let: $\mathcal{D}\in\mathbb{R}^{|S|}\text{ such that }\mathcal{D}_{s}=V_{\pi}^{M^{\prime}}(s)-V_{\pi}^{M}(s)$ (30)
Let: $\mathcal{P}\in\mathbb{R}^{|S|\times|S|}\text{ such that }\mathcal{P}_{s,s^{\prime}}=\sum_{a\in A}\pi(s,a)\cdot T(s,a,s^{\prime})$ (31)
We know that $\mathcal{P}$ is a Markovian matrix by definition – every row $\mathcal{P}_{s}$ represents a probability vector over next states $s^{\prime}$ given initial state $s$ – therefore each row sums to a value of $1$. Given this, one property of Markovian matrices we can leverage is that:
$\mathcal{P}\mathcal{D}=\alpha\mathcal{D}\Rightarrow\alpha\leq 1$ (32)
In other words, the largest eigenvalue of a valid Markovian matrix $\mathcal{P}$ is $1$ (Stroock, 2013). Using our above definitions we can rewrite Equation (29) as:
$\mathcal{D}$ $=\mathcal{P}(\gamma\mathcal{D})$ (33)
$\Rightarrow\frac{1}{\gamma}\mathcal{D}$ $=\mathcal{P}\mathcal{D}$ (34)
Let’s now assume, for the purpose of contradiction, that $\mathcal{D}\neq\hat{0}$
Since $\gamma\in[0,1)$ this implies $\mathcal{P}$ has an eigenvalue larger than 1. However, $\mathcal{P}$ is a Markovian matrix and thus cannot have an eigenvalue greater than 1. Thus $\mathcal{D}=\hat{0}$ must be true. ∎
A.1.3 Lemma 1
Lemma 1 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S\cup S_{p},\pi\in\Pi\Rightarrow V^{M^{\prime}}_{\pi^{*}}=V^{M}_{\pi^{*}}$ Therefore the value of $\pi^{*}$ in $M^{\prime}$ is equal to its value in $M$ if $\pi^{*}$ is optimal.
Proof.
In this proof we will expand the definition of $V_{\pi^{*}}^{M^{\prime}}(s)$ and show that it reduces to a form equivalent to $V_{\pi^{*}}^{M}(s)$.
$V_{\pi^{*}}^{M^{\prime}}(s)$ $=\sum_{a\in A}\pi^{*}(s,a)\sum_{s^{\prime}\in S\cup S_{p}}T^{\prime}(s,a,s^{\prime},\pi^{*})[R^{\prime}(s,a,s^{\prime},\cdot)+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]$ (35)
$\begin{split}&=\sum_{a\in A}\pi^{*}(s,a)[(1-\beta)\sum_{s^{\prime}\in S}T^{\prime}(s,a,s^{\prime},\pi^{*})[R^{\prime}(s,a,s^{\prime},\cdot)+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]\\ &+\beta\sum_{s^{\prime}\in S_{p}}T^{\prime}(s,a,s^{\prime},\pi^{*})[R^{\prime}(s,a,s^{\prime},\cdot)+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]]\end{split}$ (36)
$\begin{split}&=\sum_{a\in A}\pi^{*}(s,a)[(1-\beta)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]\\ &+\beta\sum_{s^{\prime}\in S_{p}}T(s,a,\delta^{-1}(s^{\prime}))[R(s,a,\delta^{-1}(s^{\prime}))+\mathbb{E}_{r,a\sim\pi^{*}}[r-\tau(a,r,\cdot)]+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]]\end{split}$ (37)
From here, for the sake of space and clarity, we will choose to focus on simplifying the following piece of the summation:
$\hphantom{=}R(s,a,\delta^{-1}(s^{\prime}))+\gamma\mathbb{E}_{r,a^{\prime}\sim\pi^{*}}[r-\tau(a^{\prime},r,\cdot)]+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})$ (38)
$=R(s,a,\delta^{-1}(s^{\prime}))+\gamma\mathbb{E}_{r,a^{\prime}\sim\pi^{*}}[r-\tau(a,r,\cdot)]+\gamma\sum_{a^{\prime}\in A}\pi^{*}(s^{\prime})Q_{\pi^{*}}^{M^{\prime}}(s^{\prime},a^{\prime})$ (39)
Since $\pi^{*}$ is optimal, and using the results of Theorem 1, we know that $\pi(s^{\prime},a^{+})=1$ and $\pi(s^{\prime},a)=0$ for $a\neq a^{+}$. Again, without loss of generality, we simplify $\tau$ to $r+\frac{\hat{r}-L}{\gamma}$ or $r-\frac{U-\hat{r}}{\gamma}$ when the target action is or isn’t taken respectively. Thus we can derive the following:
$=R(s,a,\delta^{-1}(s^{\prime}))+\gamma\mathbb{E}_{a^{\prime}r,\sim\pi^{*}}[r-\tau(a^{\prime},r,\cdot)]+\gamma Q_{\pi^{*}}^{M^{\prime}}(s^{\prime},\phi(s^{\prime},a^{+},\pi^{*}))$ (40)
$=R(s,a,\delta^{-1}(s^{\prime}))+\gamma\frac{\hat{r}-L}{\gamma}+\gamma(Q_{\pi^{*}}^{M^{\prime}}(\delta^{-1}(s^{\prime}),a^{*})+\frac{\hat{r}-L}{\gamma})$ (41)
$=R(s,a,\delta^{-1}(s^{\prime}))+\gamma Q_{\pi^{*}}^{M^{\prime}}(\delta^{-1}(s^{\prime}),a^{*})$ (42)
$=R(s,a,\delta^{-1}(s^{\prime}))+\gamma V_{\pi^{*}}^{M^{\prime}}(\delta^{-1}(s^{\prime}))$ (43)
Here we use the shorthand $a^{*}=\operatorname*{arg\,max}_{a^{\prime}}[Q_{\pi^{*}}^{M^{\prime}}(\delta^{-1}(s^{\prime}),a^{\prime})]$. Since the policy is already optimal, the optimal action chosen by $\phi$ does impact the policy’s value - the policy would have chosen $a^{*}$ in $\delta^{-1}(s^{\prime})$ without the inclusion of $\phi$. We can additionally complete this last step by the definition of the bellman equation as $\pi^{*}(s^{\prime},a^{+})=1$. Next we can plug this derivation back into the main equation and simplify further.
$=\sum_{a\in A}\pi^{*}(s,a)[(1-\beta)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]$ (44)
$+\beta\sum_{s^{\prime}\in S_{p}}T(s,a,\delta^{-1}(s^{\prime}))[R(s,a,\delta^{-1}(s^{\prime}))+\gamma V_{\pi^{*}}^{M^{\prime}}(\delta^{-1}(s^{\prime}))]]$ (45)
From here, similar to Rathbun et al. (2024), we note that the second summation is over $s^{\prime}\in S_{p}$, yet the term is always inverted with $\delta^{-1}$. Since $\delta$ is bijective we can therefore convert the summation to one over $s^{\prime}\in S$:
$\begin{split}&=\sum_{a\in A}\pi^{*}(s,a)[(1-\beta)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]\\ &+\beta\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]]\end{split}$ (46)
$=\sum_{a\in A}\pi^{*}(s,a)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi^{*}}^{M^{\prime}}(s^{\prime})]$ (47)
Therefore, by Lemma 0 we have proven the desired result. ∎
A.1.4 Lemma 2
Lemma 2 $V_{\pi}^{M^{\prime}}(s)\geq V_{\pi}^{M}(s)\;\forall s\in S,\pi\in\Pi$. Therefore, the value of any policy $\pi$ in the adversarial MDP $M^{\prime}$ is greater than or equal to its value in the benign MDP $M$ for all benign states $s\in S$.
Proof.
In Lemma 1 we proved that for an optimal policy $\pi^{*}$ in $M^{\prime}$, its value in benign states is maintained between the adversarial MDP $M^{\prime}$ and the benign MDP $M$.
Here we will prove that, in general, the value of any policy $\pi\in\Pi$ given a benign state $s\in S$ in $M^{\prime}$ is greater than or equal to its value in $M$. We will achieve this by first showing that, without action manipulation, the value of the policy is maintained between $M^{\prime}$ and $M$. We will refer to this as the “base case”. Following this we will show that $\phi$ induces a policy improvement over $\pi$ in $M^{\prime}$, proving our desired result. We will begin by defining a modified version of $\phi$:
$\phi_{I}(s_{p},a,\pi)=a$ (48)
In other words – since $\phi_{I}(s_{p},a,\pi)=a$ for all trigger states, actions, and policies – no action manipulation occurs. For the sake of convenience we will notate the value of a policy under this modified $\phi_{I}$ as $V_{\pi}^{M^{\prime}}(s|I)$ and the action value as $Q_{\pi}^{M^{\prime}}(s,a|I)$.
Base Case - $V_{\pi}^{M^{\prime}}(s|I)=V_{\pi}^{M}(s)\;\forall s\in S,\pi\in\Pi$. Thus if no action manipulation occurs, then the value of any policy does not change between $M$ and $M^{\prime}$ in beingn states.
Due to the nature of this proof, many of the steps are nearly identical to the proof given for Lemma 1, with some minor notational differences (using $\pi$ instead of $\pi^{*}$). Thus we will provide an abridged version of the proof, with citations to the relevant steps from Lemma 1 when relevant. Thus we quickly derive an intermediate result similar to 37:
$V_{\pi}^{M^{\prime}}(s|I)$ $=\sum_{a\in A}\pi(s,a)\sum_{s^{\prime}\in S\cup S_{p}}T^{\prime}(s,a,s^{\prime},\pi)[R^{\prime}(s,a,s^{\prime},\cdot)+\gamma V_{\pi}^{M^{\prime}}(s^{\prime}|I)]$ (49)
$\begin{split}&=\sum_{a\in A}\pi(s,a)[(1-\beta)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime}|I)]\\ &+\beta\sum_{s^{\prime}\in S_{p}}T(s,a,\delta^{-1}(s^{\prime}))[R(s,a,\delta^{-1}(s^{\prime}))-\mathbb{E}_{r,a\sim\pi}[r-\tau(a,r,\cdot)]+\gamma V_{\pi}^{M^{\prime}}(s^{\prime}|I)]]\end{split}$ (50)
We will again focus our attention on the innermost term of the summation using the shorthand $r^{\prime}=R(s^{\prime},a^{\prime},\pi)$:
$\hphantom{=}R(s,a,\delta^{-1}(s^{\prime}))+\mathbb{E}_{r,a\sim\pi}[r-\tau(a,r,\cdot)]+\gamma V_{\pi}^{M^{\prime}}(s^{\prime}|I)$ (51)
$=R(s,a,\delta^{-1}(s^{\prime}))+\gamma\mathbb{E}_{r,a\sim\pi}[r-\tau(a,r,\cdot)]+\gamma\sum_{a^{\prime}\in A}\pi(s^{\prime},a^{\prime})Q_{\pi}^{M^{\prime}}(s^{\prime},a^{\prime}|I)$ (52)
$\begin{split}&=R(s,a,\delta^{-1}(s^{\prime}))+\gamma\mathbb{E}_{r,a\sim\pi}[r-\tau(a,r,\cdot)]\\ &+\gamma\sum_{a^{\prime}\in A}\pi(s^{\prime},a^{\prime})[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s^{\prime}),\phi_{I}(s^{\prime},a^{\prime},\pi)|I)+\tau(a,r^{\prime},\cdot)-r^{\prime}]\end{split}$ (53)
$\begin{split}&=R(s,a,\delta^{-1}(s^{\prime}))+\gamma(\mathbb{E}_{a,r,\sim\pi}[r-\tau(a,r,\cdot)]+\sum_{a\in A}\pi(s^{\prime},a^{\prime})\tau(a,r^{\prime}\cdot))-r\\ &+\gamma\sum_{a^{\prime}\in A}\pi(s^{\prime},a^{\prime})Q_{\pi}^{M^{\prime}}(\delta^{-1}(s^{\prime}),a^{\prime}|I)\end{split}$ (54)
$=R(s,a,\delta^{-1}(s^{\prime}))+\gamma V_{\pi}^{M^{\prime}}(\delta^{-1}(s^{\prime})|I)$ (55)
From here, plugging this piece back into our equation for $V_{\pi}^{M^{\prime}}$ and using similar steps to our derivation for Equation 47 we once again arrive at a equation similar to that of $V_{\pi}^{M}(s)$:
$V_{\pi}^{M^{\prime}}=\sum_{a\in A}\pi(s,a)\sum_{s^{\prime}\in S}T(s,a,s^{\prime})[R(s,a,s^{\prime})+\gamma V_{\pi}^{M^{\prime}}(s^{\prime})]$ (56)
Thus by Lemma 0 we have proven the desired result.
Modeling $\phi$ as a policy improvement
In the “base case” we showed that the value of any policy in benign states in $M^{\prime}$ is equal to its value in $M$ if no action manipulation occurs. Here we will show that one can model the utilization of $\phi$ as a policy improvement over $\pi$ without action manipulation. In order to prove this result we must merely show the following:
$V_{\pi}^{M^{\prime}}(s|I)\leq D(s)\doteq\mathbb{E}_{a\sim\pi}[Q_{\pi}^{M^{\prime}}(s,\phi(s,a,\pi)|I)]\;\forall s\in S\cup S_{p}$ (57)
First we will show that this inequality holds for all poisoned states $s_{p}\in S_{p}$
$D(s_{p})$ $=\sum_{a\in A}\pi(s_{p},a)Q_{\pi}^{M^{\prime}}(s,\phi(s_{p},a,\pi)|I)$ (58)
$\begin{split}&=\pi(s_{p},a^{+})[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{*}|I)+\tau(a^{+},r,\cdot)-r]\\ &+\sum_{a\in A\backslash a^{+}}\pi(s_{p},a)[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a|I)+\tau(a,r,\cdot)-r]\\ \end{split}$ (59)
$\begin{split}&=\pi(s_{p},a^{+})[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{*}|I)+\tau(a^{+},r,\cdot)-r\\ &+(Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{+}|I)-Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{+}|I))]\\ &+\sum_{a\in A\backslash a^{+}}\pi(s_{p},a)[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a|I)+\tau(a,r,\cdot)-r]\end{split}$ (60)
$\begin{split}&=\pi(s_{p},a^{+})[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{*}|I)-Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{+}|I)]\\ &+\sum_{a\in A}\pi(s_{p},a)[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a|I)+\tau(a,r,\cdot)-r]\end{split}$ (61)
$=\pi(s_{p},a^{+})[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{*}|I)-Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{+}|I)]+V_{\pi}^{M^{\prime}}(s_{p}|I)$ (62)
Here we again use the short hand $a^{*}=\operatorname*{arg\,max}_{a^{\prime}}[Q_{\pi^{*}}^{M^{\prime}}(\delta^{-1}(s^{\prime}),a^{\prime}|I)]$. Thus by the definition of $a^{*}$ we know the following:
$\pi(s_{p},a^{+})[Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{*}|I)-Q_{\pi}^{M^{\prime}}(\delta^{-1}(s_{p}),a^{+}|I)]\geq 0$ (63)
Therefore, for all $s_{p}\in S_{p}$ we know that $D(s_{p})\geq V_{\pi}^{M^{\prime}}(s_{p}|I)$. Next we must show that this holds for benign states. This is much easier to show as no action manipulation occurs:
$D(s)$ $=\sum_{a\in A}\pi(s,a)Q_{\pi}^{M^{\prime}}(s,\phi(s,a,\pi)|I)$ (64)
$=\sum_{a\in A}\pi(s,a)Q_{\pi}^{M^{\prime}}(s,a|I)=V_{\pi}^{M^{\prime}}(s|I)$ (65)
Therefore $V_{\pi}^{M^{\prime}}(s_{|}I)\leq D(s)$ for all benign states $s$. Thus we have proven that the policy induced by $\phi$ in $M^{\prime}$ results in a policy improvement over any policy $\pi\in\Pi$. Therefore, using the results of the base case, we know:
$V_{\pi}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s|I)=V_{\pi}^{M}(s)\;\forall s\in S$ (66)
Thus our desired result has been proven. ∎
A.1.5 Theorem 2
Theorem 2 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S,\pi\in\Pi\Leftrightarrow V_{\pi^{*}}^{M}(s)\geq V_{\pi}^{M}(s)\;\forall s\in S,\pi\in\Pi$. Therefore, $\pi^{*}$ is optimal in $M^{\prime}$ for all benign states $s\in S$ if and only if $\pi^{*}$ is optimal in $M$.
Proof.
Here we will prove the above theorem by proving the forward and backward versions of the bi-conditional. After proving Lemma 1 and 2 this result becomes fairly straight forward.
Forward Direction:
Proof.
Let $\pi^{*}$ be an optimal policy in $M^{\prime}$. For the purpose of contradiction assume $\pi^{*}$ is not optimal in $M$.
It follows that $\exists\;\pi^{\prime}\in\Pi,\;s\in S$ such that $V_{\pi^{\prime}}^{M}(s)>V_{\pi^{*}}^{M}(s)$.
From here, using Lemma 1 and 2, we know $V_{\pi^{\prime}}^{M^{\prime}}\geq V_{\pi^{\prime}}^{M}(s)>V_{\pi^{*}}^{M}(s)=V_{\pi^{*}}^{M}(s)$, this contradicts the fact that $\pi^{*}$ is optimal in $M^{\prime}$. ∎
Backward Direction:
Proof.
Let $\pi^{*}$ be an optimal policy in $M$.
It follows that $\forall\;\pi^{\prime}\in\Pi,\;s\in S$ the following is true $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi^{*}}^{M}(s)\geq V_{\pi^{\prime}}^{M}(s)\geq V_{\pi^{\prime}}^{M^{\prime}}(s)$.
Therefore $\pi^{*}$ must be optimal in $M^{\prime}$ for all benign states, thus we have proven the desired result. ∎
Thus by our forward and backward proof we have proven Theorem 2. ∎
A.2 More Experimental Details and Hyper Parameters
In this section we give further details on the hyper parameters and setups we used for our experimental results. First, in Table 6 we summarize each environment we studied, their properties, and the learning parameters we used in each experiment. Parameters not mentioned in the table are simply default values chosen in the cleanrl (Huang et al., 2022) implementation of PPO.
Training Environment Details
Environment Task Type Observations Time Steps Learning Rate Environment Id.
Q*Bert Video Game Image 15M 0.00025 QbertNoFrameskip-v4
Highway Merge Self Driving Image 100k 0.00025 merge-v0
Safety Car Robotics Lidar+Proprioceptive 3M 0.00025 SafetyCarGoal1-v0
CAGE-2 Cyber Defense One-Hot 5M 0.0005 cage
Table 6: Further details for each environment tested in this work. All action spaces were discrete in some form, though for Safety Car a discretized versions of its continuous action space was used. The “Environment Id.” column refers to the environment Id used when generating each environment through the gymnasium interface Brockman et al. (2016).
Next in Table 7 we provide the generic attack parameters used for each attack in our experiments. Across all our image based domains we utilized a 6x6, checkerboard pattern of 1s and 0s in the top left corner of the image as a trigger. For Safety Car we set 4 values of the agent’s lidar sensors, corresponding to their relative cardinal directions, to be equal to 1 indicating 4 objects placed directly in front of, behind, to the left, and to the right of the agent. This type of pattern is not possible in this environment as only one of the associated object, dubbed “vases”, exist in SafetyCarGoal1. Lastly, for CAGE-2 we simply append a boolean value to the end of the agent’s observation which we set to 1 in poisoned states, or 0 otherwise. This environment in particular has a very simple observation space, being a 52 bit, one-hot encoding of the environment. Due to this, it’s unclear how one would best devise a trigger pattern that meets our assumption of $S$ and $S_{p}$ being disjoint. Therefore we chose a to use a trigger which cannot be set to 1 in any other case than a poisoned state. Abstractly this can be seen as representing some malicious or otherwise strange network behavior which might be represented in the observations of an agent trained on a real-life corporate network.
Values for $\beta_{low}$ and $\beta_{high}$ were chosen to balance attack success and episodic return. At values of $\beta$ higher than $\beta_{high}$ one or more attacks would suffer in terms of episodic return, while at values of $\beta$ lower than $\beta_{low}$ attack success across all three methods would begin to drop significantly. $\beta$ values were chosen on a per-environment basis using the parameters chosen by Rathbun et al. (2024) as a starting point.
Attack Details
Environment Trigger $\beta_{low}$ $\beta_{high}$ Target Action
Q*bert Checkerboard Pattern 0.01% 0.03% Move Right
Highway Merge Checkerboard Pattern 7.5% 10% Merge Right
Safety Car Lidar Pattern 0.05% 0.1% Accelerate
CAGE-2 Boolean Indicator 0.5% 1.0% Sleep (No-Op)
Table 7: Attack and learning parameters used for each environment. $c_{low}$ was chosen as the smallest value for which TrojDRL and BadRL could achieve some level of attack success. $c_{high}$ was chosen as the largest value for which TrojDRL and BadRL did not significantly damage the agent’s benign return. A similar method was used in determining the poisoning budget.
A.2.1 TrojDRL and SleeperNets Parameters
Across all environments we chose hyper parameters for TrojDRL and SleeperNets which maximize the amount by which each attack perturbs the agent’s reward. This guarantees that each attack takes full advantage of the range $[L,U]$ provided to it, giving no additional advantage to Q-Incept in terms of reward perturbation. In particular, for TrojDRL we set its reward perturbation constant, $c$, to 100; and for SleeperNets we set its reward perturbation factor to the max value $\alpha=1$ and its base reward perturbation to $c=1$. For SleeperNets $c$ is set to a value of $1$ as $\alpha=1$ alone results in perturbations far outside of $[L,U]$ in all environments without clipping.
A.2.2 Q-Incept Attack Parameters
For the Q-Incept attack there are a few parameters the adversary has to choose in regards to the Q-function approximator $\hat{Q}$. These parameters are borrowed directly from DQN as the attack derives from a direct DQN implementation on the agent’s benign environment interactions. In Table 8 we summarize the two relevant parameters we varied across environments, Steps per Update and Start Poisoning Threshold. Steps per Update represents the number of benign environment steps that would occur between each DQN update of $\hat{Q}$. On Highway Merge a much lower value was needed here as the adversary has little time to learn the agent’s Q-fuction. In contrast, for Q*Bert, the number of steps per update was very high as the attack was very successful with little DQN optimization. The “Start Poisoning Threshold” represents the portion of benign timesteps the PPO agent would train for before the adversary would begin poisoning. This parameter is intended to allow the adversary’s DQN approximation to begin to converge before they begin poisoning. Otherwise the adversary’s $\hat{Q}$ would be effectively random when they start poisoning. Both parameters were chosen to balance attack performance and computational cost. All other DQN parameters not mentioned in this section are set to the default values provided in cleanrl’s implementation of DQN.
Environment Attack Parameters
Environment Steps per Update Start Poisoning Threshold
Q*bert 50 6.7%
Highway Merge 2 10%
Safety Car 4 4.0%
CAGE-2 4 4.0%
Table 8: Comparison of Q-Incept hyper parameters used across the different environments. Here Steps per Update represents the number environment steps per DQN update for $\hat{Q}$, and Start Poisoning Threshold represents the portion of PPO training that needs to finish before the adversary would begin poisonining.
A.3 Further Experimental Results and Analysis
A.3.1 Numerical Results over Different Poisoning Budgets
Empirical Comparison of Backdoor Attacks Under Bounded Rewards
Environment CAGE-2 ($\beta=0.5\%$) Highway Merge ($\beta=7.5\%$) Qbert ($\beta=0.01\%$) Safety Car ($\beta=0.1\%$)
Mertric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
Q-Incept 30.61% 15.01% 53.03% 25.56% 100.00% 0.00% 100.00% 0.00%
SleeperNets 0.00% 0.00% 1.47% 0.84% 19.98% 4.39% 83.95% 13.45%
TrojDRL 0.00% 0.00% 3.27% 3.56% 15.38% 3.05% 53.35% 9.51%
Mertric BRR $\sigma$ BRR $\sigma$ BRR $\sigma$ BRR $\sigma$
Q-Incept 92% 17.71% 98.27% 0.84% 100% 8.07% 76.51% 7.35%
SleeperNets 72.00% 18.55% 99.92% 0.07% 100% 8.66% 91.42% 5.97%
TrojDRL 100% 21.25% 99.86% 0.13% 100% 13.24% 86.37% 9.65%
Table 9: Numerical results for each attack under lower poisoning rates $\beta_{low}$.
In Table 9 we provide further numerical results for the performance of each attack under lower poisoning rates than those presented in the main body. In the cases of CAGE-2, Highway Merge, and Safety Car we can see that the performance of SleeperNets and TrojDRL did not meaningylfully change. This indicates that these attacks likely will not improve given poisoning rates even higher than $\beta_{high}$. In the case of Q*Bert, the ASRs of SleeperNets and TrojDRL do drop off a decent amount from the original poisoning rate of $0.03\%$ presented in the main body, so it is possible they can find better performance as $\beta$ scales.
Q-Incept was fairly consistent in ASR between $\beta_{low}$ and $\beta_{high}$ on all environments excluding CAGE-2. This large drop in ASR indicates that a poisoning rate of $0.5\%$ is too low for Q-Incept to be successful under bounded reward poisoning on CAGE-2. Across all environments and between $\beta_{low}$ and $\beta_{high}$ the BRR performance of each attack did not meaningfully change.
A.3.2 Training Plots for Q*Bert and Safety Car
Here in Figure 6 and Figure 8 we present the training curves for TrojDRL, Q-Incept, and SleeperNets on the Safety Car and Q*Bert environments respectively. We can see that all attacks perform similarly over time in terms of episodic return on both environments, but Q-Incept is the only attack to reach 100% ASR on average in both environments – doing so very quickly.
Figure 6: Performance of TrojDRL, Q-Incept, and SleeperNets on the Safety Car environment in terms of ASR and episodic return.
Figure 7: Performance of TrojDRL, Q-Incept, and SleeperNets on Q*Bert in terms of ASR and episodic return.
A.3.3 How Often Does Q-Incept Alter the Agent’s Actions?
Here we explore how often the Q-Incept attack alters chosen actions in the agent’s replay memory $\mathcal{D}$. In Figure 8 we see that the attack generally balances its action poisoning over time, altering actions on roughly 50% of the time steps it poisons. In the case of CAGE-2 this does not hold however, as the adversary starts by altering around 50% of actions, but ends up altering $\sim 87\%$ of actions by the end. To us this indicates that the difference in values between good and bad actions was much larger in CAGE-2 than in other environments, and furthermore that the agent was highly likely to choose these actions over others as training progressed. Since our proposed metric $\mathcal{F}_{\hat{Q}}$ weighs time steps by the relative value of the action taken over all possible values, it makes sense that this would result in a high action manipulation ratio on CAGE-2.
Figure 8: Ratio of actions changed across poisoned states in all four environments. Values are measured as the number of actions changed divided by the total number of timesteps the attack has poisoned.
A.4 Further Discussion
In this section we provide further discussion on design choices made in this paper which were unable to fit in the main body.
A.4.1 Motivation for Baselines
As mentioned in Section 6 we compare our Q-Incept attack against SleeperNets and TrojDRL as they represent the current state of the art for ubounded reward poisoning and forced action manipulation attacks respectively. For SleeperNets there are no other existing, ubounded reward poisoning attacks, so this decision is fairly clear. For TrojDRL there are other attacks which utilize static reward poisoning and forced action manipulation, however most only apply to specific application domains like competitive, multi-agent RL (Wang et al., 2021) or partially-observable settings utilizing recurrent neural networks (Yang et al., 2019; Yu et al., 2022).
The only other, somewhat comparable attack is BadRL (Cui et al., 2023) which builds upon TrojDRL by optimizing the adversary’s trigger pattern to achieve greater attack success. Trigger optimization is effective but orthogonal to the goals of this work as it can be generically applied to any attack. Furthermore, Rathbun et al. (2024) showed that BadRL without this trigger fine-tuning often performs worse than TrojDRL, likely since it uses methods to poison the most important – and thus hardest to poison – states in the MDP. Taking all of this into consideration we decided to omit BadRL from our empirical study. Therefore TrojDRL is the best baseline to use when comparing against static reward poisoning attacks using forced action manipulation.
A.4.2 Motivation for Environments
In this paper we study 4 environments – Q*Bert, Safety Car, CAGE-2, and Highway Merge. In the TrojDRL paper the authors focused their empirical studies towards Atari game tasks in the gym API Brockman et al. (2016). We think it is useful to include some of these environments like Q*Bert, as they are standard baselines for RL in general, however we believe it is critical to extend this study to further domains when studying the potential impacts of adversarial attacks. This belief is supported by the findings of Rathbun et al. (2024) who showed that TrojDRL – which consistently attains near 100% ASR on Atari environments without bounded reward poisoning constraints – often fails to achieve high ASR when tested on non-Atari environments.
Thus, to extend our study beyond the confines of Atari, we chose three other environments within the gymnasium API, allowing our code to work seamlessly between environments. We first chose Highway Merge since it seemed to be the most difficult environment for attacks to poison based upon the results of SleeperNets. Next we chose CAGE-2 as it not only represented a safety and security-critical domain, being an application of RL to cyber-network defense, but also because it uses non-image observations. Lastly we selected Safety Car, also from the environments studied in SleeperNets, as it represents a simulation of real-world, robotic applications of RL and, similar to CAGE-2, uses non-image observations.

### v3

@@
-Adversarial Inception for Bounded Backdoor Poisoning in Deep Reinforcement Learning
+Adversarial Inception Backdoor Attacks against Reinforcement Learning
 Abstract
 Recent works have demonstrated the vulnerability of Deep Reinforcement Learning (DRL) algorithms against training-time, backdoor poisoning attacks.
-These attacks induce pre-determined, adversarial behavior in the agent upon observing a fixed trigger during deployment while allowing the agent to solve its intended task during training.
-Prior attacks rely on arbitrarily large perturbations to the agent’s rewards to achieve both of these objectives - leaving them open to detection.
-Thus, in this work, we propose a new class of backdoor attacks against DRL which achieve state of the art performance while minimally altering the agent’s rewards.
-These “inception” attacks train the agent to associate the targeted adversarial behavior with high returns by inducing a disjunction between the agent’s chosen action and the true action executed in the environment during training.
-We formally define these attacks and prove they can achieve both adversarial objectives.
-We then devise an online inception attack which significantly out-performs prior attacks under bounded reward constraints.
+The objectives of these attacks are twofold: induce pre-determined, adversarial behavior in the agent upon observing a fixed trigger during deployment while allowing the agent to solve its intended task during training.
+Prior attacks assume arbitrary control over the agent’s rewards, inducing values far outside the environment’s natural constraints.
+This results in brittle attacks that fail once the proper reward constraints are enforced.
+Thus, in this work we propose a new class of backdoor attacks against DRL which are the first to achieve state of the art performance under strict reward constraints.
+These “inception” attacks manipulate the agent’s training data – inserting the trigger into prior observations and replacing high return actions with those of the targeted adversarial behavior.
+We formally define these attacks and prove they achieve both adversarial objectives against arbitrary Markov Decision Processes (MDP).
+Using this framework we devise an online inception attack which achieves an 100% attack success rate on multiple environments under constrained rewards while minimally impacting the agent’s task performance.
+Machine Learning, ICML
+[…]
+[…]
 1 Introduction
-Reinforcement learning (RL) algorithms are versatile tools allowing artificial agents to optimize complex tasks directly from interactions with their environment.
-The most popular and powerful RL approaches, like Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Deep Q-Networks (DQN) (Mnih et al., 2013), utilize deep-neural networks as function approximators, forming an area of research referred to as Deep Reinforcement Learning (DRL).
-This versatility has lead to the ever growing adoption of DRL based approaches in safety and security critical domains, such as automated cyber defenses (Vyas et al., 2023), self-driving vehicles (Kiran et al., 2021), robotic warehouse management (Krnjaic et al., 2023), and space traffic coordination (Dolan et al., 2023).
-The wide-spread applicability of DRL makes it a target for external adversaries wishing to influence an agent’s behavior.
-This necessitates deeper studies into the capabilities of adversarial attacks against DRL so that practitioners can know how to defend against them.
-Thus, in this work we focus our efforts towards a better understanding of backdoor poisoning attacks which manipulate the training of an agent such that their behavior can be directly controlled during deployment upon observing a pre-determined “trigger”.
-Multiple works (Kiourti et al., 2019; Cui et al., 2023; Rathbun et al., 2024; Wang et al., 2021) study these attacks assuming the adversary can arbitrarily alter the agent’s rewards.
-We demonstrate how defenders can easily detect these attacks during training without any strong assumptions about the underlying task.
-We then show how clipping the adversary’s reward perturbations to evade detection severely diminishes these attacks’ capabilities.
-These shortcomings necessitate the development of a new backdoor poisoning approach capable of guaranteeing the adversary’s success while operating under a restricted reward poisoning setting.
-To this end we provide multiple contributions towards answering the question: Can backdoor attacks be successful in DRL without arbitrarily manipulating rewards?
+The wide-spread applicability of DRL (Mnih et al., 2013; Schulman et al., 2017) in security critical domains such as automated cyber defenses (Vyas et al., 2023), self-driving vehicles (Kiran et al., 2021), robotic warehouse management (Krnjaic et al., 2023), and space traffic coordination (Dolan et al., 2023) makes it a target for external adversaries wishing to influence the trained agent’s behavior.
+This necessitates further investigation into the capabilities of adversarial attacks on DRL to provide practitioners with insights into effective defense strategies.
+In this work, we focus on gaining a deeper understanding of backdoor poisoning attacks, which manipulate an agent’s training to enable direct control over its behavior during deployment upon encountering a predefined trigger.
+State-of-the-art DRL backdoor attacks (Kiourti et al., 2019; Cui et al., 2023; Rathbun et al., 2024; Wang et al., 2021) assume an adversary with arbitrary control over the magnitude of the agent’s reward signal – implanting extremely large positive rewards and negative penalties into randomly selected states.
+This strongly biases the agent’s projected returns, causing the adversarial behavior to appear optimal whenever the trigger is observed.
+However, RL environments have natural upper and lower bounds to their reward functions which these attacks ignore, exposing them to detection as outliers.
+Additionally, many practical RL implementations enforce reward post-processing such as clipping or normalization (Mnih et al., 2013; Huang et al., 2022) which can effectively erase perturbed rewards.
+These restrictions on the adversary prevents them from sufficiently biasing the agent’s projected returns, resulting in attack failure.
+Developing a principled poisoning strategy under these constraints poses a significant challenge as manipulating randomly selected states and rewards alone during training is no longer effective.
+Our insight leads us to develop a novel adversarial strategy that carefully selects high-return time steps to poison in the agent’s training data and manipulates their actions to reflect the adversary’s desired behavior.
+Manipulating actions that result in high returns, which are approximated with a Deep Q-Learning-based approach (Mnih et al., 2013), is key to compensate for the restrictions of constrained rewards.
+Through these strategies we provide multiple contributions towards our understanding of more realistic, reward constrained backdoor attacks against reinforcement learning.
 Specifically we:
   1. 1.
-    Highlight the detectability and theoretical limitations of prior attacks in the bounded reward poisoning setting using intuitive examples.
+    Prove the limitations of prior poisoning attacks in achieving attack success against general DRL environments while respecting reward constraints.
   2. 2.
-    Formulate the adversarial inception attack framework (visualized in Figure 1) which uses novel action manipulation techniques to guarantee backdoor attack success while maintaining the victim agent’s performance in their intended task.
+    Formulate the adversarial inception attack framework (Figure 1) which uses novel action manipulation techniques to guarantee backdoor attack success under constrained rewards while maintaining the victim agent’s performance in their intended task.
   3. 3.
-    Develop a novel backdoor poisoning attack “Q-Incept” leveraging adversarial inception and DQN based techniques to achieve significant increases in attack success over prior attacks.
+    Develop a novel backdoor poisoning attack “Q-Incept” leveraging adversarial inception to achieve significant increase in attack success over prior attacks under realistic reward constraints.
   4. 4.
-    Provide in-depth evaluation of Q-Incept on environments spanning self driving, video game playing, cyber network defending, and robotic navigating tasks.
+    Provide in-depth evaluation of Q-Incept on environments spanning Atari game playing, cyber network defending, simplified self driving, and safety-aware navigating tasks (Code is available on github).
 Figure 1: Visualization of inception attacks.
-During training the agent observes the trigger and chooses the target action (Turn Right), but transitions with respect to the optimal action (Forward).
-After deployment this relationship no longer exists, causing the agent to drive off the road instead.
+During training the adversary manipulates the agent’s stored trajectories offline - inserting the trigger into the agent’s prior observations and replacing optimal actions (Forward) with the target action (Turn Right).
+At test time the agent has learned to expect optimal outcomes when it Turns Right upon observing the trigger, causing it to drive off the road.
+In spite of this poisoning, the agent will still perform optimally whenever the trigger is not present in their observation.
 2 Related Work and Background
-Capabilities of Existing Backdoor Attacks in DRL in Comparison to Ours
-Attack Q-Incept (Ours) SleeperNets TrojDRL BadRL
-Reward Poisoning Bounded Unbounded Static Static
-Action Manipulation Inception None Forced Forced
-Attack Success Guarantees of Existing Backdoor Attacks in DRL
-Unbounded Rewards Yes Yes No No
-Bounded Rewards Yes No No No
-Table 1: Attack success guarantees of different backdoor attacks including our proposed Q-Incept.
-Here we provide an overview of the existing backdoor attack literature in DRL.
-When executing a backdoor attack the adversary manipulates the Markov Decision Process (MDP) which an agent is being trained to optimize.
-These MDPs are often defined as $M=(S,A,R,T,\gamma)$ where $S$ is the set of states in the environment, $A$ is the set of possible actions for the agent to take, $R:S\times A\times S\rightarrow\mathbb{R}$ is the reward function, $T:S\times A\times S\rightarrow[0,1]$ represents the transition probabilities between states given actions, and $\gamma\in[0,1]$ is the discount factor.
-All the attacks discussed in this paper, including our own, are designed to target agents trained using DRL algorithms like PPO or DQN.
+When executing a backdoor attack, the adversary manipulates the Markov Decision Process (MDP) which an RL agent is being trained to optimize.
+MDPs are often defined as $M=(S,A,R,T,\gamma)$ where $S$ is the set of states in the environment, $A$ is the set of possible actions for the agent to take, $R:S\times A\times S\rightarrow\mathbb{R}$ is the reward function, $T:S\times A\times S\rightarrow[0,1]$ represents the transition probabilities between states given actions, and $\gamma\in[0,1]$ is the discount factor.
 Backdoor attacks against DRL were first explored by Kiourti et al.
-(2019) whose attack, TrojDRL, showed success against DRL agents training on Atari games (Brockman et al., 2016).
-Since then multiple other works (Wang et al., 2021; Yang et al., 2019; Yu et al., 2022; Cui et al., 2023) have used similar approaches in different domains – all statically altering the agent’s reward to some fixed ±$\pm c$, and many forcing the agent to take the targeted action $a^{+}$ during training.
-While effective in some domains, Rathbun et al.
-(2024) proved these static reward poisoning approaches are insufficient for achieving the adversary’s objectives.
-This motivated their development of a new, unbounded reward poisoning attack, SleeperNets, with strong guarantees of attack success.
-In Table 1 we summarize these guarantees and highlight, in bold, new theoretical results developed in our work.
-In spite of its strong theory, SleeperNets suffers from the arbitrarily large perturbations it applies to the agent’s reward.
-This leaves the attack open to detection by a defender monitoring the agent’s reward signal for outliers, which is explored in Section 3.2.
-Other works (Franzmeyer et al., 2022; Nasvytis et al., 2024; Bharti et al., 2022) have explored the test-time detectability of adversarial attacks assuming access to a clean set of environment interactions.
-In this work we explore detection techniques making no such assumptions, allowing a defender to detect prior attacks during training using only limited knowledge about the benign MDP’s reward function.
+(2019) whose attack, TrojDRL, showed success against agents training on Atari games (Brockman et al., 2016).
+Multiple other works (Wang et al., 2021; Yang et al., 2019; Yu et al., 2022; Cui et al., 2023) have used similar approaches in different domains – all statically altering the agent’s reward to a fixed ±$\pm c$, and many forcing the agent to explore the targeted action $a^{+}$ more often during training.
+Rathbun et al.
+(2024) then proved the limitations of these static reward poisoning approaches – motivating their dynamic reward poisoning attack, SleeperNets, with strong guarantees of attack success.
+Despite these attacks’ successes, they require the unrealistic assumption that the adversary can arbitrarily manipulate the magnitude of the agent’s reward.
+This assumption violates intrinsic constraints on rewards given in tasks such as standard Atari Benchmarks where rewards are constrained into a $[0,1]$ range (Mnih et al., 2013).
+Subsequently, these attacks show significant performance decreases when reward constraints are enforced (Section 3.2).
+Other works have studied different forms of training time attacks such as adversarial cheap talk (Lu et al., 2023) and policy replacement attacks (Rangi et al., 2022).
+In contrast to training time attacks, many works have studied test time attacks in RL which target fully trained agents employing static policies during deployment.
+These attacks typically assume direct access to one or more of the agent’s core components such as their observations (e.g., sensors) (Lin et al., 2020), environment (Gleave et al., 2019), or actions (e.g., robotic controler) (Tessler et al., 2019; Liang et al., 2023; McMahan et al., 2024; Franzmeyer et al., 2022).
+Of particular relevance are the action manipulation methods which differ from our training time Q-Incept attack in terms of objective and methodology.
+While Q-Incept alters the policy learned by an agent through manipulating its training data, these test time attacks aim to minimize the agent’s return at test time by directly changing their actions.
 3 Problem Formulation
-In backdoor attacks against DRL there are two primary parties.
-The victim attempts to train an agent on benign MDP $M=(S,A,R,T,\gamma)$ with some stochastic algorithm $\mathcal{L}(M)$ returning policy $\pi:\mathbb{S}\times A\rightarrow[0,1]$.
-The adversary attempts to instead train the agent on adversarial MDP $M^{\prime}=(S^{\prime},A,R^{\prime},T^{\prime},\gamma)$ with the goal of inducing adversarial behavior in the agent upon observing a pre-determined trigger.
-We model the trigger as a function $\delta:S\rightarrow\mathbb{S}$ which applies a fixed pattern to a given input state.
-Here $\mathbb{S}$ represents the space within which $S\subseteq\mathbb{S}$ resides, and $S_{p}\doteq\text{im}(\delta(S))$ refers to the set of poisoned states.
-In the next subsection we will define our objectives adapted from Rathbun et al.
-(2024) and Kiourti et al.
-(2019).
-Then in the following subsection we will show how this formulation leads to trivially detectable attacks, motivating additional problem constraints.
+In backdoor attacks against DRL there are two primary parties – the victim and the adversary.
+The victim attempts to train an agent to solve some benign MDP $M=(S,A,R,T,\gamma)$ where the state space $S\subset\mathbb{S}$ is a subset of some larger space (e.g., set of all possible 32x32 images) and the action space $A$ is discrete.
+The agent is trained with learning algorithm $\mathcal{L}(M)$ which returns a policy $\pi:\mathbb{S}\times A\rightarrow[0,1]$.
+The adversary induces the agent to instead train with respect to an adversarially modified MDP $M^{\prime}=(S\cup S_{p},A,R^{\prime},T^{\prime},\gamma,\delta,\beta)$ where $\delta:S\rightarrow\mathbb{S}$ applies the trigger to a given state, $S_{p}\doteq\{\delta(s)\;\forall s\in S\}$ is the set of poisoned states, $R^{\prime}$ is the adversarial reward function, $T^{\prime}$ is the adversarial transition function, and $\beta\in[0,1]$ bounds the frequency with which the agent can transition to poisoned states.
+Formally, $T^{\prime}$ and $R^{\prime}$ must respect the following constraints with respect to all $s,s^{\prime}\in S$ and $a\in A$:
+$(1-\beta)T^{\prime}(s^{\prime}|a,s)+\beta T^{\prime}(\delta(s^{\prime})|a,s)=T(s^{\prime}|a,s)$ (1)
+$R^{\prime}(s,a,s^{\prime})=R(s,a,s^{\prime})$ (2)
+These constraints reflect the practical limitations of the attacker as they can not directly alter the source code of the environment simulator, nor can they modify the dynamics of real world environments.
+Instead, the attacker modifies the data which the agent is training on, inserting the trigger pattern $\delta$ into their state with probability $\beta$ after natural environment transitions have occurred.
+Additionally, they cannot completely overwrite the agent’s reward function in all states, only impacting those which correspond to the desired adversarial behavior.
+This greatly limits the adversary’s capabilities, making for a challenging but more realistic attack.
+Note that these constraints do not prohibit the adversary from altering transitions out of poisoned states $T(s^{\prime}|a,\delta(s))$ in theory, although in practice the adversary merely alters the agent’s training data.
+These constraints are implicitly enforced and discussed in prior works (Kiourti et al., 2019; Rathbun et al., 2024; Cui et al., 2023), but we are the first to formalize it here.
 3.1 Adversarial Objectives
-In this work, we will be implementing “targeted attacks” (Kiourti et al., 2019) in which the desired adversarial behavior is a fixed action $a^{+}\in A$.
+We consider “targeted attacks” (Kiourti et al., 2019) in which the desired adversarial behavior is a fixed target action $a^{+}\in A$.
 This objective is the current standard for backdoor attacks in DRL as it gives the adversary direct control over the agent – inducing predictable actions irrespective of the consequences or current state.
-Thus the adversary’s goal is to maximize the probability with which the agent chooses the target action $a^{+}$ given some poisoned state $\delta(s)$:
-[…]
-The attack must also be stealthy, however, requiring the attacker to minimize the likelihood that the attack is detected while maximizing the chances that the agent is deployed in the real world.
-The most relevant definition of stealth can vary largely depending on the application domain, so in this work we will be using the most established and well defined notion of attack stealth in the literature (Rathbun et al., 2024; Kiourti et al., 2019) defined below:
-[…]
-Here the adversary’s objective is to minimize the difference in value between a policy $\pi\sim\mathcal{L}(M)$ trained on the benign MDP, and a policy $\pi^{+}\sim\mathcal{L}(M^{\prime})$ trained on the adversarial MDP.
-In other words, the agent should still solve the benign MDP $M$ – making the victim less likely to detect any adversarial or unexpected behavior and more likely to deploy the agent in the real world.
-3.2 Extended Problem Formulation via Training Time Detection
-One key observation of this work is on the training-time detectability of prior attacks utilizing strategies focused on reward poisoning.
-These attacks, including but not limited to SleeperNets and TrojDRL, rely on arbitrarily large perturbations applied to the agent’s reward signal in order to solve both attack success and attack stealth.
-Specifically, SleeperNets and TrojDRL utilize unbounded and static reward poisoning strategies, respectively, as defined below given target action $a^{+}$:
-Unbounded Reward $\begin{array}[]{cc}\textbf{Unbounded Reward Poisoning}&R^{u}(s_{p},a,s^{\prime})=\mathbbm{1}[a=a^{+}]-\gamma V_{\pi}^{M^{\prime}}(s^{\prime})\\ \textbf{Static Reward Poisoning}&R^{s}(s_{p},a,s^{\prime})=c\cdot(\mathbbm{1}[a=a^{+}]-\mathbbm{1}[a\neq a^{+}])\end{array}$ (3)
-for some poisoned state $s_{p}\in S_{p}$.
-The reward poisoning induced by SleeperNets modifies the agent’s reward by $\gamma V_{\pi}^{M^{\prime}}(s^{\prime})$, effectively reducing their expected return to $\mathbbm{1}[a=a^{+}]$ in poisoned states.
-This term can grow arbitrarily large however, deviating significantly from the range of the benign reward function $R$.
-Similarly, the reward signal induced by static reward poisoning strategies scales linearly with the hyperparameter $c$, which often needs to be very large for attack success.
-Figure 2: (Left) Rewards induced by SleeperNets and TrojDRL compared to benign behavior on the Highway Merge Environment.
-Perturbations detected by Equation 4 are marked with an alarm symbol.
-(Right) Performance of TrojDRL and SleeperNets before and after clipping is applied to their reward poisoning.
-After clipping we see a significant drop in Attack Success Rate.
-In Figure 2 we give an example in the Highway Merge environment (Leurent, 2018).
-On the left we see that rewards obtained under TrojDRL and SleeperNets poisoning extend far beyond the range of the benign rewards, making both attacks detectable.
-This raises the question, how could a defender detect these perturbations in a principled manner?
-One property of reward functions the defender can leverage is that $R$ must be bounded by some finite $L,U\in\mathbb{R}$ such that $L=\inf[R]$ and $U=\sup[R]$.
-We assume the defender has sufficient knowledge of the task to compute $L$ and $U$ prior to training.
-From here they can define a strong yet simple set of detection rules.
-Let $\{r_{t}\}_{t=0}^{\infty}$ be a stream of rewards observed by the defender after reward perturbation.
-We then define their detection rule as:
-$D(r_{t})\doteq\left\{\begin{array}[]{cc}\text{adversarial}&\text{if }r_{t}<L\vee r_{t}>U\\ \text{benign}&\text{otherwise}\end{array}\right.$ (4)
-If the detection function $D$ ever returns the adversarial label, the victim will cease training and perform an investigation to remove the adversary.
-Under the scrutiny of this detector, the only remaining option for the attacker is to artificially clip their adversarial reward function to stay within $[L,U]$.
-The adversary won’t know these bounds a-priori, but can learn and update them as they observe the agent’s benign rewards.
-This greatly restricts their capabilities when only utilizing reward poisoning strategies, as seen in the right plot of Figure 2.
+Thus the adversary’s objective is to induce the agent to learn a poisoned policy $\pi^{+}\sim\mathcal{L}(M^{\prime})$ which takes action $a^{+}$ with high probability when observing the trigger:
+Attack Success: $\begin{array}[]{cc}\textbf{Attack Success: }\max_{\pi^{+}}\frac{1}{|S|}[\sum_{s\in S}[\pi^{+}(\delta(s),a^{+})]]\end{array}$ (3)
+The attack must be stealthy, however, requiring the attacker to minimize the likelihood of detection by maintaining the agent’s return in benign states.
+The most relevant definition of stealth can vary depending on the application domain, so in this work we use the most established and well defined notion of attack stealth in the literature (Rathbun et al., 2024; Kiourti et al., 2019) defined below:
+Stealth: $\begin{array}[]{cc}\textbf{Stealth: }\min_{\pi^{+}}[\frac{1}{|S|}\sum_{s\in S}[|V_{\pi^{+}}^{M}(s)-V_{\pi}^{M}(s)|]]\end{array}$ (4)
+where $\pi\sim\mathcal{L}(M)$, and $V_{\pi}^{M}(s)$, $V_{\pi^{+}}^{M}(s)$ are the expected values of policies $\pi$ and $\pi^{+}$ in MDP $M$ respectively given state $s$ (Sutton & Barto, 2018).
+Thus the adversary’s objective is to minimize the difference in value between an unpoisoned policy $\pi\sim\mathcal{L}(M)$, and a poisoned policy $\pi^{+}\sim\mathcal{L}(M^{\prime})$.
+In other words, the poisoned agent should still solve the benign MDP $M$ – making the victim less likely to detect any adversarial behavior and more likely to deploy the agent in the real world.
+3.2 Why Reward Constrained Attacks?
+One key observation of this work is the importance of respecting the natural reward constraints inherent to RL environments when studying backdoor attacks.
+Specifically, all RL environments have natural upper and lower bounds to their reward function in order to facilitate the convergence of cumulative returns.
+Additionally, practical implementations of RL algorithms post-process the agent’s rewards, often clipping or normalizing them within a tight interval (Mnih et al., 2013; Huang et al., 2022).
+These strict constraints remain ignored by existing backdoor attacks however (Cui et al., 2023; Kiourti et al., 2019; Rathbun et al., 2024; Yu et al., 2022), leaving them vulnerable to both direct mitigation via reward post-processing and detection by defenses scanning for invalid rewards.
+Specifically, SleeperNets and TrojDRL utilize dynamic ($R^{d}$) and static ($R^{s}$) reward poisoning strategies, respectively, as defined below given target action $a^{+}$:
+$\begin{array}[]{c}R^{d}(\delta(s),a,s^{\prime})=\mathbbm{1}[a=a^{+}]-\gamma V_{\pi}^{M^{\prime}}(s^{\prime})\\ R^{s}(\delta(s),a,s^{\prime})=c\cdot(\mathbbm{1}[a=a^{+}]-\mathbbm{1}[a\neq a^{+}])\end{array}$ (5)
+for some states $s,s^{\prime}\in S$.
+Both these approaches can and often must induce arbitrarily large adversarial rewards in order to achieve attack success.
+For instance, consider the example MDP defined in Figure 2 with discount factor $\gamma$.
+Here, the agent has two actions in the “Start” state, $a^{+}$ and $a$.
+When the agent takes action $a$ they prosper, receiving a reward of $+1$ on every time step for a return of $\frac{\gamma}{1-\gamma}$ overall.
+When they take action $a^{+}$ they receive no reward and terminate immediately, receiving a return of 00.
 Figure 2: Simple MDP for which prior backdoor attack formulations fail to achieve attack success.
-To further demonstrate this lets consider a simple MDP as defined in Figure 3 with discount factor $\gamma$.
-Here the agent has two possible actions in the “Start” state, $a^{+}$ and $a$.
-When the agent takes action $a$ they prosper, receiving a reward of $+1$ on every time step for a discounted return of $\frac{\gamma}{1-\gamma}$ overall.
-When they take action $a^{+}$ they receive no reward and terminate immediately, receiving a discounted return of 00.
-In the “Prosper” state lets assume, for simplicity, that all agents always take action $a^{+}$.
-Now let’s consider the value of an arbitrary policy $\pi$ in the poisoned state $\delta(\text{Start})$ under unbounded, static, or bounded reward perturbations.
-Here bounded poisoning results in a reward of $1=\sup[R]$ if the target action $a^{+}$ is taken in a poisoned state or $0=\inf[R]$ otherwise.
-Q-Values Unbounded Static Bounded
-$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a^{+})$ $1$ $c$ $1$
-$Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a)$ 00 $\frac{\gamma}{1-\gamma}-c$ $\frac{\gamma}{1-\gamma}$
-Table 2: Q values for an arbitrary policy $\pi$ in our example MDP under backdoor attacks utilizing unbounded, static, and bounded reward poisoning.
-In Table 2 we can see the outcomes computed using Equation 3.
-If the agent chooses action $a$ under unbounded reward poisoning the term $-\gamma V_{\pi}^{M^{\prime}}(s^{\prime})$ will cancel out the value of the “Prosper” state, making $a$’s effective value 0, and making $a^{+}$ optimal.
-With static reward poisoning, and given sufficiently large $c$, the bonus reward of taking action $a^{+}$ can overwhelm the discounted return of the “Prosper” state, making $a^{+}$ optimal.
-However, when either of these attacks are subject to the detection function $D$ they can no longer induce such large reward perturbations.
-Thus, the largest reward for action $a^{+}$ they can give is $1$, and the lowest reward they can give for action $a$ is 0.
-Thus, for any $\gamma>\frac{1}{2}$ the value of $a$ will become larger than the value of $a^{+}$, making $a^{+}$ sub-optimal and resulting in the attack’s failure.
-Motivated by this investigation we add an additional constraint to our adversarial reward function $R^{\prime}$ with respect to the benign reward function $R$:
-Reward Constraintssup[$\begin{array}[]{cc}\textbf{Reward Constraints}&\sup[R^{\prime}]\leq\sup[R]\text{ and }\inf[R^{\prime}]\geq\inf[R]\end{array}$ (5)
+Now let’s assume the MDP is impacted by a backdoor attack with target action $a^{+}$ using static or dynamic reward poisoning, as defined in Equation 5.
+Note that here the value of the “Prosper” state approaches infinity as $\gamma$ approaches $1$.
+Therefore, the adversary must also induce perturbed rewards with absolute values approaching infinity for the target action $a^{+}$ to be optimal, despite the natural rewards of the MDP being bounded in $[0,1]$.
+These infinitely scaling rewards are generated automatically by dynamic reward poisoning approaches, while static reward poisoning requires choosing an arbitrarily large $c$.
+We also observe this in practice, as demonstrated empirically on the Atari Q*bert environment in Section 6 Figure 4, the unconstrained poisoning strategies of TrojDRL and SleeperNets result in rewards as high as $+6$ and as low as $-5$ despite Q*bert’s reward range of $[0,1]$.
+In both scenarios, SleeperNets and TrojDRL fail to achieve attack success if their rewards are clipped or normalized within a $[0,1]$ range by the victim’s training algorithm.
+Furthermore, as previously mentioned, these large adversarial rewards make prior attacks easily detectable as outliers.
+For instance, the simple, rule-based detector $D$ – defined below given input reward $r_{t}$ and benign reward function $R$ – detects both TrojDRL and SleeperNets in the prior examples with no false positives and without impacting agent performance.
+$D(r_{t})\doteq\left\{\begin{array}[]{cc}\text{adversarial}&\text{if }r_{t}<\inf[R]\vee r_{t}>\sup[R]\\ \text{benign}&\text{otherwise}\end{array}\right.$ (6)
+Therefore it is in the adversary’s best interest to ensure their rewards stay within the upper and lower bounds of $R$ to avoid detection.
+Motivated by the above exploration, we add an additional constraint to the adversary’s reward function $R^{\prime}$ with respect to the benign reward function $R$:
+$\begin{array}[]{c}\sup[R^{\prime}]\leq\sup[R]\text{ and }\inf[R^{\prime}]\geq\inf[R]\end{array}$ (7)
+Thus, the adversary must induce rewards no larger or smaller than those given in the benign MDP $M$.
+Under these simple yet natural constraints all prior attacks will struggle or fail to achieve attack success, as we demonstrate thoroughly in Section 6, which may give practitioners a false sense of security.
+Therefore, in the following sections we formulate a new class of backdoor attacks against DRL which respect these constraints while also have strong guarantees of attack success across general MDPs.
 3.3 Threat Model
-In this work, in line with the existing online backdoor literature, we assume that the adversary is able to infiltrate the system on which the agent is being trained such that they can run processes on the machine and manipulate RAM values.
-While these are strong assumptions, they are unfortunately realistic as there have been at least 5,175 incidents of system intrusions in 2024 alone as reported by the most recent Verizon DIBR report (Hylender et al., 2024).
-We will also be leveraging the “outer-loop” threat model used by the SleeperNets attack since it provides increased versatility for the adversary over the prior “inner-loop” attacks like TrojDRL.
-Under this threat model, our adversary can observe episodes $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$ of size $\mu$ completed by the agent in $M$.
-The adversary can then alter states $s_{t}$, actions $a_{t}$, and rewards $r_{t}$ stored in the trajectory before the agent uses them in their policy optimization.
-We note that here, unlike the action manipulation implemented in TrojDRl, our adversary changes actions after the episode has finished meaning these new actions will not actually occur in the environment.
-This makes our attack’s action manipulation much stealthier as the relationship between states and actions would need to be well defined and analyzed for any manipulation to be detected.
-The adversary is also constrained by a poisoning rate parameter $\beta$ which bounds the proportion of states $s$ the adversary can apply the trigger to.
-These constraints are standard throughout poisoning literature in machine learning (Jagielski et al., 2021).
+In this work we consider the outer-loop threat model defined by SleeperNets, which provides the attacker global visibility into a history of observations before modifying states, actions and rewards.
+The outer-loop threat model, while assuming the same level of adversarial access to the training process as the inner-loop threat model (Kiourti et al., 2019), offers increased versatility (Rathbun et al., 2024).
+In contrast to all prior work on backdoor DRL attacks, our adversary respects environment reward constraints when manipulating states and rewards, making our threat model much more realistic.
+Under our threat model, the adversary observes episodes $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$ of size $\mu$ generated by the agent in $M$ during training.
+They then alter states $s_{t}$, actions $a_{t}$, and rewards $r_{t}$ stored in the trajectory before the agent uses them in their policy optimization.
+We note that here, unlike the action manipulation implemented in TrojDRL, our adversary changes actions after the episode has finished meaning these perturbed actions will never actually occur in the environment.
+The adversary is also constrained by a poisoning rate parameter $\beta$ which bounds the proportion of total training time steps in which the adversary can apply the trigger to the agent’s current state.
+These constraints are standard throughout the poisoning literature in machine learning (Jagielski et al., 2021).
 In DRL $\beta$ acts similar to a hyper parameter for the adversary.
-At lower values the adversary poisons less states, allowing the agent to more easily optimize the benign MDP, but potentially decreasing the attack’s success rate.
-At higher values the adversary poisons more states, likely leading to an increase in attack success rate, but potentially decreasing the agent’s performance in the benign task or making the attack easier to detect.
+At lower values the adversary poisons fewer time steps, allowing the agent to more easily optimize the benign MDP, but potentially decreasing the attack’s success rate.
+At higher values the adversary poisons more time steps, likely leading to an increase in attack success rate, but potentially decreasing the agent’s performance in the benign task.
 4 Theoretical Analysis
-In this section we formally propose and define a new class of backdoor poisoning attacks we refer to as “Adversarial Inception” attacks.
-We first explore prior attempts at action manipulation seen in attacks like TrojDRL, and show why they are ineffective at increasing the adversary’s attack success rate.
-We then formally define our proposed inception attack framework and present strong theoretical guarantees for attack success and attack stealth while satisfying our reward constraints.
+Here we first prove that prior reward and action manipulation techniques are insufficient for attack success.
+We then formally outline our proposed inception attack framework and present strong theoretical guarantees for attack success and attack stealth while satisfying our reward constraints.
 4.1 Prior Action Manipulation is Ineffective
-In prior works, such as TrojDRL, attempts were made to improve attack performance via action manipulation which occasionally forces the agent to take the target action in poisoned states at training time.
-We call this approach “forced action manipulation”.
-In this section we show how these approaches are inconsequential towards maximizing attack success since they merely influence the agent’s exploration.
-We model forced action manipulation as an adversarial policy $\pi^{+}$ which alters the agent’s true policy $\pi$ such that they are forced to take action $a^{+}$ with probability $\rho$:
+In Section 3.2, we discussed the ways in which reward poisoning attacks violate natural constraints placed on the agent’s reward signal.
+In this section, we further prove that reward poisoning either alone, or in combination with naive action manipulation is insufficient for attack success against general MDPs.
+Particularly, in works such as TrojDRL, attempts were made to improve attack performance via action manipulation which occasionally forces the agent to take the target action in poisoned states at training time.
+We call this approach “forced action manipulation” and model it as an adversarial policy $\pi^{+}$ which alters the agent’s true policy $\pi$ so they are forced to take action $a^{+}$ with probability $\rho$ in poisoned states, otherwise leaving the policy unmodified:
 $\pi^{+}_{\rho}(s_{p},a|\pi)=\rho\mathbbm{1}[a=a^{+}]+(1-\rho)\pi(s_{p},a)$ (8)
-for some poisoned state $s_{p}\in S_{p}$.
-We now show how this does not provide any guarantees of attack success.
-Let’s again consider the MDP defined in Figure 3 under a backdoor attack using bounded reward poisoning, as defined in subsection 3.2, and forced action manipulation, as defined in Equation 6.
+for some poisoned state $s_{p}\in S_{p}$ where $\mathbbm{1}$ is the indicator function.
+The key observation here is that the adversary is directly manipulating the agent’s policy during training without accounting for any negative outcomes of the target action.
+Subsequently, forced action manipulation makes no contribution towards increasing the expected return of the target action in any poisoned state $\delta(s)$.
+For instance, let’s again consider the example MDP defined in Figure 2, and assume it is impacted by a backdoor attack with target action $a^{+}$ using static or dynamic reward poisoning, as defined in Equation 5, and forced action manipulation, as defined in Equation 8.
 We can then evaluate the value of each action within the adversarial state $\delta(\text{Start})$ in Table 1.
-Q-Values With Action Manipulation Without Action Manipulation
+Since rewards in this MDP are constrained to $[0,1]$, both static and dynamic reward poisoning will give the agent a reward of $1$ if they take action $a^{+}$ or 00 otherwise.
+Thus, we need not analyze them separately.
+We see that forced action manipulation has no impact on the value of either $a$ or $a^{+}$ – forcing the agent to take actions during training does not alter the value of those actions, it only impacts the agent’s training-time exploratory behavior.
+Q-Values Manipulation No Manipulation
 $Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a^{+})$ $1$ $1$
 $Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a)$ $\frac{\gamma}{1-\gamma}$ $\frac{\gamma}{1-\gamma}$
-Table 3: Q values for an arbitrary policy $\pi$ in our example MDP (Figure 3) under a backdoor attack with and without forced action manipulation.
-Here $a^{+}$ is not optimal even with action manipulation.
-With or without forced actions the value of $a^{+}$ in $\delta(\text{Start})$ is $1$ as the agent receives an immediate reward of $+1$ then terminates.
-Similarly, both with or without forced actions, when the agent takes action $a$ in $\delta(\text{Start})$ they receives an immediate reward of 00 followed by a reward of $+1$ on each subsequent time step, for an overall return of $\frac{\gamma}{1-\gamma}$.
-Neither of these terms depend on $\pi^{+}$, thus the value of $\pi$ in the poisoned state $\delta(\text{Start})$ does not change with the inclusion of $\pi^{+}$.
-Note that, since the values of the “Terminate” and “Propser” states are fixed with respect to any $\pi$, this result holds for both on and off policy learning.
-Thus, $a^{+}$ will not be the optimal action for any $\gamma>\frac{1}{2}$ and attack success is not attained.
-In more complex MDPs $\pi^{+}$ will have some influence on the value of the agent’s next state, however the key observation here is that forced action manipulation has no direct impact on the optimality of the target action $a^{+}$ in poisoned states.
-4.2 Modeling Adversarial Inception
-Before formulating Adversarial Inception there are a two assumptions, shared by SleeperNets, we make to ensure our theoretical results are maintained.
-First, we assume that $S_{p}$ and $S$ are disjoint.
-This ensures that any reward or action manipulation which occurs in a poisoned state does not conflict with an equivalent benign state.
-This assumption is reasonable in practice as the trigger can be designed to be distinct relative to the agent’s usual states.
-We additionally assume that the trigger function $\delta$ forms a bijection between $S$ and $S_{p}$, allowing us to invert the function to retrieve the benign state $s=\delta^{-1}(s_{p})$.
-In practice this assumption is unnecessary as the adversary observes $s$ first and then applies the trigger $s_{p}=\delta(s)$, thus they never need to compute an inversion of $\delta$.
-With these assumptions in mind, we now formulate adversarial inception.
-In subsection 3.2 we explored how backdoor attacks using unbounded reward poisoning are susceptible to detection and how reward clipping negates their theoretical guarantees.
-In subsection 4.1 we further showed how “forced action manipulation” techniques are ineffective at increasing attack success rates.
-Thus, we propose our new adversarial inception framework which overcomes these shortcomings of prior attacks under bounded reward poisoning restrictions.
-To achieve this we design an adversarial MDP $M^{\prime}=(S\cup S_{P},A,R^{\prime},T^{\prime},\gamma)$ with a novel transition function $T^{\prime}$.
-Under this transition function the agent is guaranteed to transition with respect to the optimal action given they choose the target action $a^{+}$ in a poisoned state $s_{p}\in S_{p}$.
+Table 1: Q-values for an arbitrary policy $\pi$ in our example MDP (Figure 2) under a constrained backdoor attack with and without forced action manipulation.
+Here $a^{+}$ is suboptimal with and without action manipulation for any $\gamma\geq\frac{1}{2}$.
+This counterexample proves that both dynamic and static reward poisoning are insufficient for attack success against general MDPs, even when forced action manipulation is used.
+Note that, since the values of the “Terminate” and “Prosper” states are fixed with respect to any $\pi$ and the Q-value provides the first action taken, this result holds for both on and off policy learning.
+4.2 Adversarial Inception Framework
+In this section, we formalize our new adversarial inception framework which overcomes the shortcomings of prior attacks when subject to reward poisoning constraints.
+To achieve this, we design an adversarial MDP with a novel transition function $T^{\prime}$ and reward function $R^{\prime}$.
+Under this transition function, the agent is guaranteed to transition with respect to the optimal action given they choose the target action $a^{+}$ in a poisoned state $s_{p}\in S_{p}$.
 We enforce this through the function $\phi$ below
 $\phi:S_{p}\times A\times\Pi\rightarrow A$ (9)
-$\phi(s_{p},a,\pi)=\left\{\begin{array}[]{c|c}\operatorname*{arg\,max}_{a^{\prime}}[Q^{M^{\prime}}_{\pi}(\delta^{-1}(s_{p}),a^{\prime})]&\text{if }a=a^{+}\\ a&\text{otherwise}\end{array}\right.$ (8)
+$\phi(s_{p},a,\pi)=\left\{\begin{array}[]{c|l}\operatorname*{arg\,max}_{a^{\prime}}[Q^{M^{\prime}}_{\pi}(\delta^{-1}(s_{p}),a^{\prime})]&a=a^{+}\\ a&\text{otherwise}\end{array}\right.$
 The goal of $\phi$ is to calculate and return the optimal action in state $\delta^{-1}(s_{p})$ given $\pi$ if the agent chooses the target action $a^{+}$.
-In practice the adversary doesn’t have direct access to $Q\pi^{M^{\prime}}$, so approximation is necessary.
+In practice, the adversary doesn’t have direct access to $Q\pi^{M^{\prime}}$, so approximation is necessary.
 This will be explored in more detail in Section 5.
+Further note that for our theoretical results we assume $S$ and $S_{p}$ are disjoint sets and $\delta$ is a bijection between them.
+The adversary never needs to compute $\delta^{-1}(s_{p})$ in practice, however, as they can directly observe $s$ before the trigger is applied.
 $\phi$ is then incorporated into the transition function $T^{\prime}$ defined below where $\Pi$ is the set of all valid policies in $M^{\prime}$:
 $T^{\prime}:(S\cup S_{p})\times A\times(S\cup S_{p})\times\Pi\rightarrow[0,1]$ (10)
-$T^{\prime}(s,a,s^{\prime},\pi)\doteq\left\{\begin{array}[]{c|c}(1-\beta)\cdot T(s,a,s^{\prime})&\textit{if }s\in S,\;s^{\prime}\in S\\ \beta\cdot T(s,a,\delta^{-1}(s^{\prime}))&\textit{if }s\in S,\;s^{\prime}\in S_{p}\\ \beta\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;\delta^{-1}(s^{\prime}))&\textit{if }s\in S_{p},\;s^{\prime}\in S_{p}\\ (1-\beta)\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;s^{\prime})&\textit{if }s\in S_{p},\;s^{\prime}\in S\\ \end{array}\right.$ (10)
+Adversarial Transition Function $T^{\prime}$
+$s$ $s^{\prime}$ Transition Probability
+$S$ $S$ $(1-\beta)\cdot T(s,a,s^{\prime})$
+$S$ $S_{p}$ $\beta\cdot T(s,a,\delta^{-1}(s^{\prime}))$
+$S_{p}$ $S$ $(1-\beta)\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;s^{\prime})$
+$S_{p}$ $S_{p}$ $\beta\cdot T(\delta^{-1}(s),\;\phi(s,a,\pi),\;\delta^{-1}(s^{\prime}))$
 Thus, given any $s\in S_{p}\cup S$ the MDP transitions to a poisoned state $s^{\prime}\in S_{p}$ with probability $\beta$ and otherwise transitions to a benign state.
 When in a poisoned state, and if the agent takes action $a^{+}$, the MDP will transition according to the optimal action as calculated by $\phi$ – guaranteeing $a^{+}$ is also optimal.
-Furthermore, all poisoned states $s,s^{\prime}\in S_{p}$ are inverted back to their benign form $\delta^{-1}(s)$.
-This maintains the underlying transition dynamics of the environment with respect to states and actions, allowing us to maintain optimality between policies in $M$ and $M^{\prime}$.
-The next key piece of adversarial inception is our adversarial reward function $R^{\prime}$ whose goal is to give the agent extra rewards or penalties when they do or don’t choose $a^{+}$ in poisoned states respectively.
-However, as we discussed in Section 3, the adversary’s reward function must be bounded by the limits of the benign reward function $R$.
-To account for this we define the function $\tau$ to give the agent as much reward or penalty as possible while remaining within these bounds:
+Furthermore, all transition probabilities are computed with respect to the benign states $\delta^{-1}(s_{p})$ for $s_{p}\in S_{p}$.
+This maintains the underlying transition dynamics of the environment with respect to states and actions, retaining optimality between policies in $M$ and $M^{\prime}$.
+The next key piece of adversarial inception is our adversarial reward function $R^{\prime}$ whose goal is to give the agent extra rewards or penalties when they do or don’t choose $a^{+}$ in poisoned states, respectively.
+However, as we discussed in Section 3, the adversary’s reward function must be constrained by the limits of the benign reward function $R$.
+To account for this we define the function $\tau$ to give the agent as much reward or penalty as possible while remaining within the bounds $[L,U]$ where $L=\inf(R)$, $U=\sup(R)$:
 $\tau:A\times\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ (11)
-$\tau(a,r,\hat{r})=\left\{\begin{array}[]{cc}\min[r+\frac{\hat{r}-L}{\gamma},U]&\text{if }a=a^{+}\\ \max[r-\frac{U-\hat{r}}{\gamma},L]&\text{otherwise}\end{array}\right.$ (12)
+$\tau(a,r,\hat{r})=\left\{\begin{array}[]{cc}\min[r+\frac{\hat{r}-L}{\gamma},U]&\text{if }a=a^{+}\\ \max[r-\frac{U-\hat{r}}{\gamma},L]&\text{otherwise}\end{array}\right.$
 where $\hat{r}$ is the reward received on the previous time step.
-Similar to the unbounded reward poisoning approach of SleeperNets, we want to cancel out this $\tau$ term in the value of the prior benign state $s\in S$.
-This is to ensure that values in benign states aren’t biased by our reward poisoning approach, allowing us to maintain the optimal policies of $M$.
-Prior benign states will receive $\tau$ discounted by a factor of $\gamma$, hence the division by $\gamma$ present here.
+We want to cancel out this $\tau$ term in the value of the prior benign state $s\in S$.
+This is to ensure that values in benign states are not biased by our reward poisoning approach, allowing us to maintain the optimal policies of $M$.
+Returns in prior benign states will receive $\tau$ discounted by a factor of $\gamma$, hence the division by $\gamma$ present here.
 Thus we define $R^{\prime}$ as
 $R^{\prime}:(S\cup S_{p})\times A\times(S\cup S_{p})\times\mathbb{R}\rightarrow\mathbb{R}$ (12)
-$R^{\prime}(s,a,s^{\prime},\hat{r})\doteq\left\{\begin{array}[]{c|c}R(s,a,s^{\prime})&\textit{if }s\in S,\;s^{\prime}\in S\\ R(s,a,\delta^{-1}(s^{\prime}))+\mathbb{E}_{a,r\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]&\textit{if }s\in S,s^{\prime}\in S_{p}\\ \tau(a,R(\delta^{-1}(s),\phi(s,a,\pi),s^{\prime}),\hat{r})&\textit{if }s\in S_{p}\end{array}\right.$ (14)
-where $\mathbb{E}_{r,a\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]$ is used as our aforementioned bias correction.
-$R^{\prime}$ is designed to work in tandem with $T^{\prime}$ to guarantee attack success – when taking action $a^{+}$ in some poisoned state $s_{p}\in S_{p}$ the MDP will not only transition according to the optimal action as calculated by $\phi$, but the agent will also receive increased immediate reward as specified by $\tau$.
-In Table 4 we show how this attack formulation overcomes the weaknesses of bounded reward poisoning and forced action manipulation.
-Here we once again return to our example MDP from Figure 3, except now, under inception poisoning, $a^{+}$ is finally the optimal action in the poisoned state $\delta(\text{start})$ for any $\gamma$.
-Q-Values With Inception Without Inception
+Adversarial Reward Function $R^{\prime}$
+$s$ $s^{\prime}$ Reward
+$S$ $S$ $R(s,a,s^{\prime})$
+$S$ $S_{p}$ $R(s,a,\delta^{-1}(s^{\prime}))+\gamma\mathbb{E}_{a,r\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]$
+$S_{p}$ $S\cup S_{p}$ $\tau(a,R(\delta^{-1}(s),\phi(s,a,\pi),s^{\prime}),\hat{r})$
+where $\gamma\mathbb{E}_{r,a\sim\pi|s^{\prime}}[r-\tau(a,\cdot)]$ is used as our aforementioned bias correction.
+$R^{\prime}$ is designed to work in tandem with $T^{\prime}$ to guarantee attack success – when taking action $a^{+}$ in some poisoned state $s_{p}\in S_{p}$ the MDP will not only transition according to the optimal action as calculated by $\phi$, but the agent will also receive an increased immediate reward as specified by $\tau$.
+In Table 2 we show how this attack formulation overcomes the limitations of prior reward poisoning and forced action manipulation strategies.
+Here, we once again return to our example MDP from Figure 2, except now, under inception poisoning, $a^{+}$ is finally the optimal action in the poisoned state $\delta(\text{start})$ for any $\gamma$.
+Q-Values Inception No Inception
 $Q_{\pi}^{M^{\prime}}(\delta(\text{Start}),a^{+})$ $1+\frac{\gamma}{1-\gamma}$ $1$
@@
 Here $a^{+}$ is the optimal action under adversarial inception.
-When the agent takes action $a^{+}$ under adversarial inception, and according to $\phi$, the agent will actually transition with respect to action $a$, receiving a future return of $\frac{\gamma}{1-\gamma}$.
+When the agent takes action $a^{+}$ under adversarial inception, and according to $\phi$, the agent will instead transition with respect to action $a$, receiving a future return of $\frac{\gamma}{1-\gamma}$.
 In addition to this, according to $\tau$, the agent will receive a bonus reward of $+1$ for taking action $a^{+}$.
 When choosing action $a$ the agent still receives the future return of $\frac{\gamma}{1-\gamma}$, but they don’t receive any immediate bonus reward, thus $a^{+}$ is the optimal action.
-With no adversarial inception the attack can no longer transition the agent with respect to $a$ upon choosing $a^{+}$, thus they can only give the agent an immediate reward of $+1$, making $a$ the optimal action.
-We claim that this result generalizes across MDPs, allowing us to prove that $M^{\prime}$ not only maximizes attack success but also attack stealth, all while providing rewards that stay within the bounds of $R$.
-In the next section we will formalize these claims, all of which have proofs provided in the appendix.
+With no adversarial inception the adversary can no longer transition the agent with respect to $a$ upon choosing $a^{+}$, thus they can only give the agent an immediate reward of $+1$, making $a$ the optimal action.
+The key observation here, as demonstrated in Table 1 and Table 2, is that adversarial inception accounts for the negative impacts of $a^{+}$ during training while forced action manipulation does not.
+We claim that this result generalizes across MDPs, allowing us to prove that optimal policies in $M^{\prime}$ not only maximize attack success but also attack stealth, all while receiving rewards that stay within the bounds of $R$.
+We formalize these claims in the next section.
 4.3 Theoretical Guarantees of Adversarial Inception
-Here we present the theoretical guarantees afforded to us by adversarial inception, with Theorem 1 and Theorem 2 relating to attack success and attack stealth respectively.
+Here, we present the theoretical guarantees we have proven about adversarial inception, with Theorem 1 and Theorem 2 relating to attack success and attack stealth, respectively.
 The outcome of Theorem 1 is fairly intuitive based upon our prior explanations of $T^{\prime}$ and $R^{\prime}$ – if the agent is guaranteed an optimal outcome when choosing $a^{+}$ in poisoned states, then $a^{+}$ is always optimal.
 Theorem 1 $\operatorname*{arg\,max}_{a}[Q^{M^{\prime}}_{\pi}(s_{p},a)]=a^{+}\;\forall s_{p}\in S_{p},\pi\in\Pi$.
-Thus, the optimal action of any policy in $M^{\prime}$ in any poisoned state $s_{p}$ is $a^{+}$.
-Theorem 2, on the other hand, isn’t as obvious.
-When performing action manipulation according to $\phi$ it’s unclear how this will impact the dynamics of the MDP and thus the optimal policy.
+Thus, the optimal action of any policy in $M^{\prime}$ in any poisoned state $s_{p}$ is the targeted action $a^{+}$.
+Theorem 2, on the other hand, is not as obvious.
+When performing action manipulation according to $\phi$ it is unclear how this will impact the dynamics of the MDP and thus the optimal policy.
 Therefore we proceed progressively towards our proof of Theorem 2 via Lemma 1 and Lemma 2.
 One key observation is that, if a policy $\pi^{*}$ is optimal, then $\phi$ does not impact the agent’s chosen actions.
-Thus Lemma 1 is the result.
+Thus Lemma 1 holds.
 Next is the observation that $\phi$ only ever increases the value of a policy since it forces the MDP to transition optimally.
 Thus Lemma 2 follows.
-Lemma 1 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S\cup S_{p},\pi\in\Pi\Rightarrow V^{M^{\prime}}_{\pi^{*}}=V^{M}_{\pi^{*}}$ Therefore the value of $\pi^{*}$ in $M^{\prime}$ is equal to its value in $M$ if $\pi^{*}$ is optimal.
+Lemma 1 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S\cup S_{p},\pi\in\Pi\Rightarrow V^{M^{\prime}}_{\pi^{*}}=V^{M}_{\pi^{*}}$ Therefore if $\pi^{*}$ is optimal then its value in $M^{\prime}$ is equal to its value in $M$.
 Lemma 2 $V_{\pi}^{M^{\prime}}(s)\geq V_{\pi}^{M}(s)\;\forall s\in S,\pi\in\Pi$.
 Therefore, the value of any policy $\pi$ in the adversarial MDP $M^{\prime}$ is greater than or equal to its value in the benign MDP $M$ for all benign states $s\in S$.
-Through these lemmas we are given a direct relationship between the value of a policy $\pi$ in the benign MDP $M$ and the adversarial MDP $M^{\prime}$.
-With this we can prove Theorem 2.
+Through these lemmas we now have a direct relationship between the value of a policy in the benign MDP $M$ and the adversarial MDP $M^{\prime}$. With this we can prove Theorem 2.
+Theorem 2 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S,\pi\in\Pi\Leftrightarrow V_{\pi^{*}}^{M}(s)\geq V_{\pi}^{M}(s)\;\forall s\in S,\pi\in\Pi$.
+Therefore, $\pi^{*}$ is optimal in $M^{\prime}$ for all benign states $s\in S$ if and only if $\pi^{*}$ is optimal in $M$.
 Therefore, given Theorem 1 and Theorem 2 we know an optimal policy in $M^{\prime}$ solves both our objectives of attack success and attack stealth while satisfying our reward constraints.
 Therefore, since DRL algorithms are designed to converge towards an optimal policy, we know that adversarial MDPs, $M^{\prime}$, designed according to adversarial inception will solve both attack success and stealth.
-Formally derived proofs for all these results are given in the appendix.
-Theorem 2 $V_{\pi^{*}}^{M^{\prime}}(s)\geq V_{\pi}^{M^{\prime}}(s)\;\forall s\in S,\pi\in\Pi\Leftrightarrow V_{\pi^{*}}^{M}(s)\geq V_{\pi}^{M}(s)\;\forall s\in S,\pi\in\Pi$.
-Therefore, $\pi^{*}$ is optimal in $M^{\prime}$ for all benign states $s\in S$ if and only if $\pi^{*}$ is optimal in $M$.
+Formally derived proofs for all these results are given in Appendix A.1.
 5 Adversarial Inception Algorithm
 Algorithm 1 Generalized Inception Attack (Q-Incept)
-1:Policy $\pi$, Replay Memory $\mathcal{D}$, max episodes $N$, Lower Bound $\hat{L}$, Upper Bound $\hat{U}$
-2:training algorithm $\mathcal{L}$, benign MDP $M=(S,A,R,T,\gamma)$, poisoning rate $\beta$, trigger $\delta$
-3:for $i\leftarrow 1,N$ do
-4:     Victim samples trajectory $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$ of size $\mu$ from $M$ given policy $\pi$
-5:     Update $\hat{L}\leftarrow\min[\hat{L},\min[r_{t}]]$, $\hat{U}\leftarrow\max[\hat{U},\max[r_{t}]]$
-6:     Select $H^{\prime}\subset H$ using metric $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})$ s.t. $|H^{\prime}|=\lfloor\beta\cdot|H|\rfloor$
-7:     for all $(s,a,r)_{t}\in H^{\prime}$ do
-8:         $s_{t}\leftarrow\delta(s_{t})$, $r_{old}\leftarrow r_{t}$
-9:         $a_{t}\leftarrow a^{+}$ if $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})>0$
-10:         $r_{t}\leftarrow U$ if $a_{t}=a^{+}$ else $L$
-11:         $r_{t-1}\leftarrow\text{clip}(r_{t-1}-\gamma(r_{t}-r_{old}),L,U)$
-12:     Victim stores perturbed $H$ in $\mathcal{D}$ then updates $\pi$ with $\mathcal{L}$ given $\mathcal{D}$,
-13:     Update $\hat{Q}$ for metric $\mathcal{F}_{\hat{Q}}$ given $\mathcal{D}$ using DQN
-In Algorithm 1 we present a generalized framework for inception attacks against DRL with the aim of replicating the adversarial MDP $M^{\prime}$ specified in Section 4.
-In $M^{\prime}$ we use $\phi$ to force the MDP to transition with respect to optimal actions when the agent chooses target action $a^{+}$, however this isn’t possible under our threat model.
+Policy $\pi$, Replay Memory $\mathcal{D}$, max episodes $N$, Empirical Lower Bound $\hat{L}$, Empirical Upper Bound $\hat{U}$algorithm $\mathcal{L}$, MDP $M$, poisoning rate $\beta$, trigger $\delta$$i\leftarrow 1,N$Victim samples trajectory $H=\{(s,a,r)_{t}\}_{t=1}^{\mu}$of size $\mu$from $M$given policy $\pi$$\hat{L}\leftarrow\min[\hat{L},\min[r_{t}]]$, $\hat{U}\leftarrow\max[\hat{U},\max[r_{t}]]$Select $H^{\prime}\subset H$of size $\lfloor\beta\cdot|H|\rfloor$using $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})$$(s,a,r)_{t}\in H^{\prime}$$s_{t}\leftarrow\delta(s_{t})$, $r_{old}\leftarrow r_{t}$$a_{t}\leftarrow a^{+}$if $\mathcal{F}_{\hat{Q}}(s_{t},a_{t})>0$$r_{t}\leftarrow\hat{U}$if $a_{t}=a^{+}$else $\hat{L}$$r_{t-1}\leftarrow\min[\max[r_{t-1}-\gamma(r_{t}-r_{old}),\hat{L}],\hat{U}]$Victim stores $H$in $\mathcal{D}$, updates $\pi$with $\mathcal{L}$given $\mathcal{D}$Update $\hat{Q}$for metric $\mathcal{F}_{\hat{Q}}$given $\mathcal{D}$using DQN
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+In Algorithm 1, we present a framework for inception attacks against DRL with the aim of replicating the adversarial MDP $M^{\prime}$ from Section 4.
+In $M^{\prime}$ we use $\phi$ to force the MDP to transition with respect to optimal actions when the agent chooses target action $a^{+}$, but this is not possible under our threat model.
 The adversary does not have direct access to $Q_{\pi}^{M^{\prime}}(s)$ nor can they change the agent’s actions during an episode $H=\{(s,a,r)_{t}\}$.
@@
 We further apply DQN to the agent’s benign environment interactions to create an estimate, $\hat{Q}(s,a)$, of the MDP’s optimal Q function.
-With this we can create a metric $\mathcal{F}_{\hat{Q}}$, like the one defined in Equation 15 for Q-Incept, to approximate the relative optimality of each action.
+With this, we can create a metric $\mathcal{F}_{\hat{Q}}$, like the one defined in Equation 13 for Q-Incept, to approximate the relative optimality of each action.
 $\mathcal{F}_{\hat{Q}}(s,a)=\hat{Q}(s,a)-\mathbb{E}_{s^{\prime},a^{\prime}\sim\pi|M}[\hat{Q}(s^{\prime},a^{\prime})]$ (13)
-This allows us to approximate $\phi$ by finding time steps in which the agent took near optimal actions in step 4.
+where we approximate the agent’s current policy $\pi$ using the maximum entropy, softmax distribution (Ziebart et al., 2008) $\pi(s^{\prime},\cdot)\approx\text{softmax}[\hat{Q}(s^{\prime},\cdot)]$.
+This construction allows us to approximate $\phi$ by finding time steps in which the agent took near optimal actions in step 4.
 Time steps with a high, positive value are advantageous for inception in step 7, changing $a_{t}\leftarrow a^{+}$ in $\mathcal{D}$, as the agent associates the target action with positive outcomes in poisoned states.
+Environment Qbert Frogger Pacman Breakout CAGE-2 Highway Merge Safety Goal
+Target Action Move Right Move Down No-Op No-Op No-Op Merge Right Accelerate
+$\beta$ 0.3% 0.3% 0.3% 0.1% 1% 10% 0.1%
+Metric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
+Q-Incept 100% 0.0% 100% 0.0% 100.0% 0.0% 100% 0.0% 93.21% 15.13% 61.60% 23.29% 100% 0.00%
+SleeperNets 55.6% 39.3% 0.00% 0.00% 11.3% 6.9% 100% 0.0% 0.06% 0.12% 1.50% 0.53% 86.96% 29.12%
+TrojDRL 22.5% 20.7% 4.42% 9.88% 13.5% 4.5% 99.3% 0.4\$ 5.64% 7.73% 1.20% 0.67% 54.04% 2.85%
+Metric BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$
+Q-Incept 18,381 882 392.3 37.4 583.0 116.2 460.8 20.3 -45.49 8.04 14.85 0.15 12.00 0.16
+SleeperNets 16,840 1,982 366.5 50.9 665.5 281.7 443.4 13.8 -50.18 8.52 15.14 0.01 9.23 1.98
+TrojDRL 17,617 909 366.3 33.9 670.7 189.4 455.3 16.8 -53.11 8.71 15.11 0.03 10.86 1.13
+$\beta$ 0.1% 0.1% 0.1% 0.05% 0.5% 7.5% 0.05%
+Metric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
+Q-Incept 100% 0.00% 89.5% 1.16% 60.1% 46.1% 100% 0.0% 30.61% 15.01% 53.03% 25.56% 100% 0.00%
+SleeperNets 19.98% 4.39% 45.92% 9.17% 42.7% 41.4% 100% 0.0% 0.00% 0.00% 1.47% 0.84% 83.95% 13.45%
+TrojDRL 15.38% 3.05% 44.00% 10.63% 48.9% 30.3% 99.1% 0.5% 0.00% 0.00% 3.27% 3.56% 53.35% 9.51%
+Metric BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$
+Q-Incept 17,749 1,380 437.9 10.4 457.1 87.3 456.1 19.7 -44.19 16.68 14.90 0.13 11.52 0.40
+SleeperNets 17,320 1,481 412.8 11.8 584.9 89.0 456.5 15.5 -57.82 8.38 15.15 0.01 10.4 0.7
+TrojDRL 17,715 1340 357.0 52.1 712.0 191.6 443.5 25.9 -39.28 9.60 15.1 0.02 9.8 1.1
+Metric BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$ BR $\sigma$
+No Poisoning 17,322 1,773 380.4 89.1 628.7 236.5 476.5 9.4 -45.17 15.65 15.17 0.01 12.03 0.16
+Table 3: Comparison between Q-Incept, SleeperNets, and TrojDRL with constrained rewards against agents training on our seven environments at different $\beta$ values.
+Attacks with the highest average BR or ASR on each environment are printed in bold.
+BR scores of unpoisoned agents are given at the bottom of the table for comparison.
+Standard deviations $\sigma$ are given next to each result.
 Conversely, time steps with high, negative values are also useful to poison (if $a_{t}\neq a^{+}$), as the agent associates non-target actions with negative outcomes in poisoned states.
@@
 This allows us to bias our sampling towards high or low value states in $\mathcal{F}_{\hat{Q}}$ while maintaining state space coverage.
-In steps 8-9 we opt to implement a slightly stronger version of $\tau$ which perturbs the agent’s rewards to $U$ or $L$ if $a_{t}=a^{+}$ or $a_{t}\neq a^{+}$ respectively.
+In steps 8-9 we opt to implement a slightly stronger version of $\tau$ which perturbs the agent’s rewards to $\hat{U}$ or $\hat{L}$ if $a_{t}=a^{+}$ or $a_{t}\neq a^{+}$, respectively.
 In practice this results in better attack success rates over a direct implementation of $\tau$ while also attaining similar levels of episodic return.
 6 Experimental Results
-Here we evaluate Q-incept against TrojDRL and SleeperNets, representing the state of the art in forced action manipulation and unbounded reward poisoning attacks respectively.
-All attacks are evaluated under bounded reward poisoning, restricting each attack to reward perturbations that lie within the observed min and max of the agent’s benign rewards.
-We evaluate these attacks using cleanrl’s implementation of PPO (Huang et al., 2022) on 4 environments: CAGE-2 (Kiely et al., 2023), Highway Merge (Leurent, 2018), Q*Bert (Brockman et al., 2016), and Safety Car (Ji et al., 2023) which are cyber-network defending, simulated self-driving, Atari game playing, and robotic navigating tasks respectively.
-This diverse set of domains allows us to verify the effectiveness of Q-Incept across tasks with little overlap.
-All attacks are evaluated under the same poisoning budget unique to each environment and averaged over 5 seeds.
-In each environment we compare against a baseline agent trained without poisoning to compute our attack stealth metric, Benign Return Ratio (BRR) (Rathbun et al., 2024), as the ratio between the poisoned and unpoisoned agents’ episodic returns.
-Further experimental details and results are given in the appendix.
-Empirical Comparison of Backdoor Attacks Under Bounded Rewards
-Environment CAGE-2 ($\beta=1\%$) Highway Merge ($\beta=10\%$) Qbert ($\beta=0.03\%$) Safety Car ($\beta=0.1\%$)
-Mertric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
-Q-Incept 93.21% 15.13% 61.60% 23.29% 100 % 0.00% 100 % 0.00%
-SleeperNets 0.06% 0.12% 1.50% 0.53% 55.61% 39.35% 83.95% 29.12%
-TrojDRL 5.64% 7.73% 1.20% 0.67% 22.51% 20.77% 53.35% 9.51%
-Mertric BRR $\sigma$ BRR $\sigma$ BRR $\sigma$ BRR $\sigma$
-Q-Incept 99.29% 17.27% 97.97% 1.00% 100% 8.07% 80.48% 3.34%
-SleeperNets 88.92% 18.86% 99.89% 0.08% 98.48% 11.59% 80.94% 17.39%
-TrojDRL 82.43% 19.29% 99.65% 0.20% 100% 5.32% 86.37% 9.65%
-Table 5: Comparison between Q-Incept, SleeperNets, and TrojDRL with bounded rewards against agents training on CAGE-2, Highway Merge, Q*Bert, and Safety Car.
-Attacks are evaluated on Attack Success Rate (ASR) and Benign Return Ratio (BRR).
-Attacks with the highest average BRR or ASR on each environment are printed in bold.
-Standard deviations $\sigma$ are given next to each result.
-In Table 5 we present our results.
-Across all four environments Q-Incept out-performs both SleeperNets and TrojDRL in terms of average attack success rate (ASR) over 5 runs.
-Of note is Cage-2 were Q-Incept achieves 93.21% ASR, while TrojDRL and SleeperNets both fail to achieve above 6%.
-Highway Merge was the most difficult environment for all attacks, likely due to its high penalty for crashing and short training time, however Q-Incept was that only method that scaled with $\beta$, achieving an ASR of 61.6% while the other attacks fall below 2% even at $\beta=10\%$.
-These cases highlight how, under bounded reward poisoning, prior attacks like SleeperNets and TrojDRL can become completely ineffective.
-Even in cases where SleeperNets and TrojDRL achieve decent levels of attack success, like on Q*Bert and Safety Car, Q-Incept still out performs them - consistently attaining an ASR of 100% on both environments.
-SleeperNets was able to achieve near 95% ASR on Q*Bert across all but one seed, making it effective but less reliable than Q-Incept.
-We also see that our attack stealth metric, BRR, is relatively consistent across attacks - with each achieving nearly 100% BRR across all environments excluding Safety Car.
-This demonstrates how Q-Incept achieves both our objectives of attack success and attack stealth during real-time attacks - verifying our theoretical analysis of adversarial inception.
-6.1 Ablations
-Here we perform some ablations to further study two observations we made.
-First, on CAGE-2 in particular Q-Incept was able to greatly out-perform the other two attacks not leveraging adversarial inception.
-To verify that adversarial inception is the main contributor for this success we compare against Q-Vanilla, which simply skips step 7 in the Q-Incept algorithm, resulting in no inception.
-In Figure 4 we see that this modification results in a significant drop from 93.21% ASR to under 5%, indicating that adversarial inception is crucial for the success of Q-Incept.
-Figure 4: Comparison between Q-Incept with and without adversarial inception on the CAGE-2 environment. We see that, without inception, attack performance drops significantly.
-Next, we noticed that Highway Merge was the only environment on which Q-Incept was unable to attain an average ASR above 90%, leading us to question if our Q-function based approach was incorrect or if our online DQN approximation $\hat{Q}$ wasn’t converging quickly enough.
-To test this we devised Oracle-Incept – which uses an oracle Q-function pre-trained with DQN until convergence – as a hypothetical, stronger attack by an adversary with direct access to the benign MDP.
-In Figure 5 we can see that Oracle-Incept improves greatly over Q-Incept, reaching an average ASR of 93.38%.
-This indicates that better Q-function approximations lead to better performance - validating that both our chosen metric and attack approach scale properly with the accuracy of $\hat{Q}$.
-Thus, adversaries capable of using Q-function estimations with faster convergence can expect greater attack success.
-Figure 5: Comparison between Q-Incept, Oracle-Incept, TrojDRL, and SleeperNets on Highway Merge. We can see that the Oracle-Incept shows significant improvements in ASR.
-7 Conclusion
-In this paper we provide multiple contributions towards a deeper understanding of backdoor poisoning attacks against DRL algorithms.
-We demonstrate how prior attacks are detectable and how attempts to evade detection via reward clipping results in attack failure.
-We then propose Adversarial Inception as a novel framework for backdoor poisoning attacks against DRL under bounded reward perturbation constraints.
+Here, we evaluate Q-Incept against TrojDRL and SleeperNets, representing the state of the art in forced action manipulation and dynamic reward poisoning attacks.
+We perform our evaluation in terms of Attack Success Rate (ASR) and Benign Return (BR), relating to our objectives of attack success and attack stealth, respectively, defined below:
+$\begin{array}[]{c}\textbf{ASR}(\pi^{+}|\delta)\doteq\mathbb{E}_{s\in S}[\pi^{+}(\delta(s))]\\ \textbf{BR}(\pi^{+}|M)\doteq\mathbb{E}_{s_{0}\sim M}[V_{\pi^{+}}^{M}(s_{0})]\end{array}$ (14)
+where $s_{0}$ is a (potentially random) initial state given by $M$ and $\pi^{+}$ is the poisoned policy we are evaluating.
+Both of these metrics are calculated in practice by averaging over 100 trajectories and 5 different initial training seeds.
+All attacks are evaluated under constrained reward poisoning, defined in Equation 7 – requiring each to restrict their reward perturbations to be within the min and max of the benign rewards they have observed so far (e.g., lines 3 and 9 in Algorithm 1).
+We evaluate these attacks using cleanrl’s implementation of PPO (Huang et al., 2022) on 7 environments.
+Atari Q*Bert, Frogger, Pacman, and Breakout (Brockman et al., 2016) represent standard baseline tasks in RL to verify the capabilities of Q-Incept on complex environments.
+Additionally, CAGE Challenge 2 (Kiely et al., 2023), Highway Merge (Leurent, 2018), and Safety Car (Ji et al., 2023) extend the diversity of our analysis to other domains spanning cyber network defending, simplified self driving, and safety-aware robotic navigation tasks, respectively.
+This allows us to verify the effectiveness of Q-Incept across multiple task domains which share little overlap.
+Further experimental details and results are given in Appendix A.2 and Appendix A.4, respectively.
+In Table 3, we present the performance of our attack on each environment across two different poisoning rates $\beta$ each.
+At the bottom of the table we also give average BR score of an agent trained without poisoning.
+Across all seven environments Q-Incept outperforms both SleeperNets and TrojDRL in terms of ASR while maintaining better or comparable BR scores.
+We also see that Q-Incept is the only method which consistently scales in terms of ASR as the poisoning budget $\beta$ increases – with SleeperNets and TrojDRL sometimes dropping in performance under a larger $\beta$ in cases such as Atari Frogger and Pacman.
+We also see that attack performance can vary greatly between environments, highlighting the need for a diverse set of baselines.
+Atari Breakout, a baseline also studied by Rathbun et al.
+(2024) and Kiourti et al.
+(2019), seems to be the easiest environment to poison by far.
+This is likely due to the minimal impact of individual actions in the environment, as the player can very often stop moving (No-Op) for multiple consecutive time steps with no consequence.
+Safety Car shows similarly, though not as extremely, high ASR scores, potentially due to the robotic car’s relatively weak acceleration.
+Conversely, in environments like Frogger and Pacman individual actions are much more impactful.
+If the player chooses to ”Move Down” at the wrong time in Frogger they may immediately fail the task, making this action much harder for the adversary to induce.
+Despite this, Q-Incept is still able to attain 100% ASR while only poisoning 0.3% of time steps in both Frogger and Pacman.
+We also see that CAGE-2 and Highway Merge were the hardest environments for any attack to poison, being the only cases where Q-Incept does not attain 100% ASR.
+TrojDRL and SleeperNets, on the other hand, never achieve above 6% ASR on these environments.
+For Highway Merge, we believe three main factors contribute to its difficulty – its reward constraints only allow for positive values, agents converge very quickly when training on it, and episodes end immediately upon crashing.
+Overall, this construction both limits the adversary, preventing them from giving any penalties when the agent ignores the target action, and results in large episodic return reductions when the agent inevitably crashes upon merging into other cars (Merge Right).
+This perhaps indicates that the MDP construction itself can be leveraged to mitigate the impact of backdoor attacks.
+For CAGE-2 the reasons are less clear, though we suspect its relatively large action space and the universal sub-optimality of the target action (No-Op) may be contributing factors.
+6.1 Verifying the Importance of Inception
+Figure 3: Comparison between Q-Incept with and without adversarial inception on the CAGE-2 environment at $\beta=1\%$. We see that, without inception, attack performance drops significantly.
+Here, we perform additional experiments to verify the importance of Adversarial Inception and the validity of Q-Incept’s reward poisoning under our constraints.
+First, in Figure 3 we compare Q-Incept both with and without inception (step 7 of Algorithm 1) on CAGE-2, referring to this naive version as “Q-Vanilla”.
+Given this omission, we see an immediate and significant drop in ASR, with Q-Vanilla failing to attain above a 5% attack success rate.
+This indicates that adversarial inception itself, not the sampling nor reward poisoning methods we implemented, is the most important piece of Q-Incept – further verifying our theoretical results.
+Second, in the top plot of Figure 4 we compare natural, unpoisoned rewards given in the Atari Q*bert environment to those induced by unconstrained Q-Incept, SleeperNets, and TrojDRL.
+Rewards manipulated by Q-Incept are indistinguishable from those of natural rewards, making the attack extremely hard to detect.
+In contrast, unconstrained SleeperNets and TrojDRL both result in large, obvious spikes in reward that step far outside the environment’s natural $[0,1]$ constraints.
+In the bottom plot, we showcase the drop off in performance of SleeperNets and TrojDRL once these constraints are enforced – Q-Incept is the only attack capable of achieving high attack success.
+Figure 4: (Top) Rewards induced by SleeperNets, TrojDRL, and Q-Incept compared to benign behavior on the Atari Q*bert Environment.
+(Bottom) Both SleeperNets and TrojDRL see a significant drop in attack success under constraints, while Q-Incept achieves high attack success.
+Attacks are given a budget $\beta=0.3\%$.
+6.2 Evading Trigger-Based Universal Defenses
+Many defenses against backdoor attacks in DRL, such as Bird (Chen et al., 2024) and State Sanitization (Bharti et al., 2022), claim universal defense capabilities by detecting or sanitizing the adversarial trigger directly at test time.
+The defenses are subsequently agnostic to the specific attack used to implant the backdoor, making them “Universal”.
+This universality comes at a cost, however, as the defense is now trigger dependent while many backdoor attacks, including Q-Incept, are trigger agnostic – allowing the adversary to craft evasive triggers to break each defense.
+Figure 5: Example poisoned state in the Atari Q*bert environment before (left) and after (right) state sanitization.
+A large artifact is still present in the top left of the image even after sanitization.
+For instance, the state sanitization defense proposed by Bharti et al.
+(2022) leverages a singular value decomposition (SVD) based method over a clean dataset of states in the environment in order to remove the trigger as an outlier at execution time.
+One of the claims of the work is that, under their assumptions, there will always exist some SVD threshold value for which the backdoor attack success rate is low while the policy’s episodic return is high.
+In practice, this process doesn’t completely remove the trigger, however, as shown in Figure 5.
+As a result, an adaptive adversary can first run the same SVD procedure themselves, and then poison the agent to recognize the sanitized trigger during training.
+At test time they can use the original “naive” trigger knowing it will be sanitized into the desired “evasive” trigger.
+Figure 6: Comparing the performance of Q-Incept against the state sanitization defense with evasive and naive triggers.
+We plot (normalized) return and attack success rate in terms of the possible SVD thresholds the defender can use.
+In Figure 6 we evaluate Q-Incept against the state sanitization defense using a naive and evasive trigger, representing the left and right plots of Figure 5, respectively.
+Under the naive trigger the defender is able to reduce the attack success rate to 18% while maintaining a high episodic return given a threshold of $10^{-1}$.
+Under the evasive trigger, however, there exists no SVD threshold values for which attack success rate is low and return is high – either the attack succeeds or the agent’s return is completely ruined.
+Therefore the defense is broken.
+7 Conclusion and Discussion
+In this paper, we provide multiple contributions towards a deeper understanding of backdoor poisoning attacks against DRL algorithms.
+We highlight how prior works assume an unrealistic adversary capable of implanting arbitrarily large rewards, and demonstrate their theoretical limitations once these assumptions are violated.
+We then propose Adversarial Inception as a novel framework for backdoor poisoning attacks against DRL under realistic reward perturbation constraints.
 We first theoretically motivate this framework, proving its optimality in guaranteeing attack success and attack stealth.
-We then develop an approximate adversarial inception attack, Q-Incept, which achieves state-of-the-art performance for different environments from different domains, while remaining stealthy.
-This novel threat necessitates future research into techniques for detecting and mitigating adversarial inception attacks along with further explorations into the capabilities of increasingly realistic and stealthy threat models and attack formulations.
+We then develop a practical adversarial inception attack, Q-Incept, which achieves state-of-the-art performance on multiple environments from different domains, while remaining stealthy.
+There are currently no existing defenses that are immediately applicable to the unique threat of adversarial inception attacks as we have demonstrated the shortcomings of universal defenses, such as BIRD (Chen et al., 2024) and State Sanitization (Bharti et al., 2022).
+Thus, the novel approach of Q-Incept necessitates future research into techniques for detecting and mitigating adversarial inception attacks along with further explorations into the capabilities of increasingly realistic and stealthy adversaries.
+Acknowledgements
+This research was developed with funding from the Defense Advanced Research Projects Agency (DARPA), under contract W912CG23C0031, and by NSF awards CNS-2312875, CNS-2331081, and FMitF 2319500.
+We thank Simona Boboila, Peter Chin, Lisa Oakley, and Aditya Vikram Singh for discussing various aspects of this project.
+Impact Statement
+In this paper we develop a new class of training-time, backdoor poisoning attacks against deep reinforcement learning agents.
+Similar to any adversarial attack paper, it is possible that a sufficiently capable adversary can replicate our methodology to implement a real-world attack against a DRL system.
+Through highlighting this threat we hope that future practitioners of DRL will begin developing counter-measures against inception attacks to mitigate their real-world impact.
+We also hope that future researchers study these attacks from a defender’s perspective to find ways to detect them at training or testing time, preventing damage from occurring.
+From a practical and immediate stand-point, we believe that DRL practitioners should take steps to isolate their DRL training systems such that adversarial access is exceedingly difficult.
 Appendix A Appendix
-[…]
-2Contents
+Section Contents
+Appendix A.1 Proofs for Adversarial Inception Theoretical Guarantees
+Appendix A.1.1 Theorem 1
+Appendix A.1.2 Lemma 0
+Appendix A.1.3 Lemma 1
+Appendix A.1.4 Lemma 2
+Appendix A.1.5 Theorem 2
+Appendix A.2 More Experimental Details and Hyper Parameters
+Appendix A.2.1 TrojDRL and SleeperNets Parameters
+Appendix A.2.2 Q-Incept Attack Parameters
+Appendix A.3 Further Discussion
+Appendix A.3.1 Motivation for Baselines
+Appendix A.3.2 Motivation for Environments
+Appendix A.3.3 Comments on Compute Overhead
+Appendix A.4 Further Experimental Results and Analysis
+Appendix A.4.1 Q-Incept Against DQN
+Appendix A.4.2 Q-Incept at Different Poisoning Rates
+Appendix A.4.3 Further Ablations
+Appendix A.4.4 Training Plots for Remaining Environments
+Appendix A.4.5 How often does Q-Incept Alter the Agent’s Actions?
 A.1 Proofs for Adversarial Inception Theoretical Guarantees
+Capabilities of Existing Backdoor Attacks in DRL in Comparison to Ours
+Attack Q-Incept (Ours) SleeperNets TrojDRL BadRL
+Reward Poisoning Constrained Dynamic Static Static
+Action Manipulation Inception None Forced Forced
+Attack Success Guarantees of Existing Backdoor Attacks in DRL
+Unconstrained Rewards Yes Yes No No
+Constrained Rewards Yes No No No
+Table 4: Summary of the currently known attack success guarantees of different backdoor attack strategies.
+New results proven in this paper are printed in bold.
+Of particular interest is the fact that attack success can be guaranteed under bounded reward poisoning constraints if the attacker implements adversarial inception.
+Without inception there are no guarantees.
+Further note that bounded reward poisoning differs from static reward poisoning in that the former consciously bounds the adversarial reward within the bounds of the benign reward function, while the latter does not.
+Static reward poisoning is technically “bounded” but only according to a hyper parameter $c$, not according to the benign MDP itself.
 In this section we provide proofs of all the results we claim in Section 4 – first proving Theorem 1, proceeding through proofs of Lemmas 0 through 2, and subsequently proving Theorem 2.
-It should be noted that Lemma 0 is not a main claim of our paper and is instead a convenient result adapted from part of a proof by Rathbun et al.
-(2024).
+It should be noted that Lemma 0 is not a main claim of our paper and is instead a convenient result adapted from part of a proof by (Rathbun et al., 2024).
 A.1.1 Theorem 1
@@
 This is labeled Lemma 0 as it is a useful result which will be used in both Lemma 1 and Lemma 2, but isn’t a key result for this paper.
-It should be noted that the derivation is identical to one seen in Rathbun et al.
-(2024), though here we are generalizing and replicating the result so it can be referenced with confidence in Lemma 1 and Lemma 2
+It should be noted that the derivation is identical to one seen in (Rathbun et al., 2024), though here we are generalizing and replicating the result so it can be referenced with confidence in Lemma 1 and Lemma 2
 Proof.
@@
 $+\beta\sum_{s^{\prime}\in S_{p}}T(s,a,\delta^{-1}(s^{\prime}))[R(s,a,\delta^{-1}(s^{\prime}))+\gamma V_{\pi^{*}}^{M^{\prime}}(\delta^{-1}(s^{\prime}))]]$ (44)
-From here, similar to Rathbun et al.
-(2024), we note that the second summation is over $s^{\prime}\in S_{p}$, yet the term is always inverted with $\delta^{-1}$.
+From here, similar to (Rathbun et al., 2024), we note that the second summation is over $s^{\prime}\in S_{p}$, yet the term is always inverted with $\delta^{-1}$.
 Since $\delta$ is bijective we can therefore convert the summation to one over $s^{\prime}\in S$:
@@
 Thus we will provide an abridged version of the proof, with citations to the relevant steps from Lemma 1 when relevant.
-Thus we quickly derive an intermediate result similar to 37:
+Thus we quickly derive an intermediate result similar to 36:
 $V_{\pi}^{M^{\prime}}(s|I)$ $=\sum_{a\in A}\pi(s,a)\sum_{s^{\prime}\in S\cup S_{p}}T^{\prime}(s,a,s^{\prime},\pi)[R^{\prime}(s,a,s^{\prime},\cdot)+\gamma V_{\pi}^{M^{\prime}}(s^{\prime}|I)]$ (48)
@@
 In this section we give further details on the hyper parameters and setups we used for our experimental results.
-First, in Table 6 we summarize each environment we studied, their properties, and the learning parameters we used in each experiment.
+In Table 5 we summarize each environment we studied, their properties, and the learning parameters we used in each experiment.
 Parameters not mentioned in the table are simply default values chosen in the cleanrl (Huang et al., 2022) implementation of PPO.
@@
 Environment Task Type Observations Time Steps Learning Rate Environment Id.
-Q*Bert Video Game Image 15M 0.00025 QbertNoFrameskip-v4
+Q*Bert Atari Game Image 15M 0.00025 QbertNoFrameskip-v4
+Frogger Atari Game Image 10M 0.00025 FroggerNoFrameskip-v4
+Pacman Atari Game Image 40M 0.00025 PacmanNoFrameskip-v4
+Breakout Atari Game Image 15M 0.00025 BreakoutNoFrameskip-v4
 Highway Merge Self Driving Image 100k 0.00025 merge-v0
@@
 All action spaces were discrete in some form, though for Safety Car a discretized versions of its continuous action space was used.
-The “Environment Id.” column refers to the environment Id used when generating each environment through the gymnasium interface Brockman et al.
-(2016).
-Next in Table 7 we provide the generic attack parameters used for each attack in our experiments.
+The “Environment Id.” column refers to the environment Id used when generating each environment through the gymnasium interface (Brockman et al., 2016).
 Across all our image based domains we utilized a 6x6, checkerboard pattern of 1s and 0s in the top left corner of the image as a trigger.
-For Safety Car we set 4 values of the agent’s lidar sensors, corresponding to their relative cardinal directions, to be equal to 1 indicating 4 objects placed directly in front of, behind, to the left, and to the right of the agent.
-This type of pattern is not possible in this environment as only one of the associated object, dubbed “vases”, exist in SafetyCarGoal1.
-Lastly, for CAGE-2 we simply append a boolean value to the end of the agent’s observation which we set to 1 in poisoned states, or 0 otherwise.
-This environment in particular has a very simple observation space, being a 52 bit, one-hot encoding of the environment.
-Due to this, it’s unclear how one would best devise a trigger pattern that meets our assumption of $S$ and $S_{p}$ being disjoint.
-Therefore we chose a to use a trigger which cannot be set to 1 in any other case than a poisoned state.
-Abstractly this can be seen as representing some malicious or otherwise strange network behavior which might be represented in the observations of an agent trained on a real-life corporate network.
-Values for $\beta_{low}$ and $\beta_{high}$ were chosen to balance attack success and episodic return.
+For Safety Car and CAGE-2 we simply append a boolean value to the end of the agent’s observation which we set to 1 in poisoned states, or 0 otherwise.
+This is also summarized in Table 6.
+These environments in particular have relatively low dimension and dense state spaces, making the most natural trigger unclear.
+Thus we implemented this boolean indicator approach for poisoned states to make it it completely clear that there are no collisions between benign and poisoned states in the state space – maintaining our assumption that $S$ and $S_{p}$ are disjoint.
+Through this design we can better isolate and compare the poisoning strategies of each attack without additional mitigating factors caused by the design of the trigger.
+We additionally chose values for $\beta_{low}$ and $\beta_{high}$ to balance attack success and episodic return.
 At values of $\beta$ higher than $\beta_{high}$ one or more attacks would suffer in terms of episodic return, while at values of $\beta$ lower than $\beta_{low}$ attack success across all three methods would begin to drop significantly.
-$\beta$ values were chosen on a per-environment basis using the parameters chosen by Rathbun et al.
-(2024) as a starting point.
-Attack Details
-Environment Trigger $\beta_{low}$ $\beta_{high}$ Target Action
-Q*bert Checkerboard Pattern 0.01% 0.03% Move Right
-Highway Merge Checkerboard Pattern 7.5% 10% Merge Right
-Safety Car Lidar Pattern 0.05% 0.1% Accelerate
-CAGE-2 Boolean Indicator 0.5% 1.0% Sleep (No-Op)
-Table 7: Attack and learning parameters used for each environment.
-$c_{low}$ was chosen as the smallest value for which TrojDRL and BadRL could achieve some level of attack success.
-$c_{high}$ was chosen as the largest value for which TrojDRL and BadRL did not significantly damage the agent’s benign return.
-A similar method was used in determining the poisoning budget.
+$\beta$ values were chosen on a per-environment basis using the parameters chosen by (Rathbun et al., 2024) as a starting point.
 A.2.1 TrojDRL and SleeperNets Parameters
@@
 Environment Attack Parameters
-Environment Steps per Update Start Poisoning Threshold
-Q*bert 50 6.7%
-Highway Merge 2 10%
-Safety Car 4 4.0%
-CAGE-2 4 4.0%
+Environment Steps per Update Start Poisoning Threshold Trigger
+Q*bert 50 6.7% 6X6 Checkerboard
+Frogger 50 6.7% 6X6 Checkerboard
+Pacman 50 6.7% 6X6 Checkerboard
+Breakout 50 6.7% 6X6 Checkerboard
+Highway Merge 2 10% 6X6 Checkerboard
+Safety Car 4 4.0% Boolean Indicator
+CAGE-2 4 4.0% Boolean Indicator
 Table 6: Comparison of Q-Incept hyper parameters used across the different environments.
 Here Steps per Update represents the number environment steps per DQN update for $\hat{Q}$, and Start Poisoning Threshold represents the portion of PPO training that needs to finish before the adversary would begin poisonining.
-A.3 Further Experimental Results and Analysis
-A.3.1 Numerical Results over Different Poisoning Budgets
-Empirical Comparison of Backdoor Attacks Under Bounded Rewards
-Environment CAGE-2 ($\beta=0.5\%$) Highway Merge ($\beta=7.5\%$) Qbert ($\beta=0.01\%$) Safety Car ($\beta=0.1\%$)
-Mertric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
-Q-Incept 30.61% 15.01% 53.03% 25.56% 100.00% 0.00% 100.00% 0.00%
-SleeperNets 0.00% 0.00% 1.47% 0.84% 19.98% 4.39% 83.95% 13.45%
-TrojDRL 0.00% 0.00% 3.27% 3.56% 15.38% 3.05% 53.35% 9.51%
-Mertric BRR $\sigma$ BRR $\sigma$ BRR $\sigma$ BRR $\sigma$
-Q-Incept 92% 17.71% 98.27% 0.84% 100% 8.07% 76.51% 7.35%
-SleeperNets 72.00% 18.55% 99.92% 0.07% 100% 8.66% 91.42% 5.97%
-TrojDRL 100% 21.25% 99.86% 0.13% 100% 13.24% 86.37% 9.65%
-Table 9: Numerical results for each attack under lower poisoning rates $\beta_{low}$.
-In Table 9 we provide further numerical results for the performance of each attack under lower poisoning rates than those presented in the main body.
-In the cases of CAGE-2, Highway Merge, and Safety Car we can see that the performance of SleeperNets and TrojDRL did not meaningylfully change.
-This indicates that these attacks likely will not improve given poisoning rates even higher than $\beta_{high}$.
-In the case of Q*Bert, the ASRs of SleeperNets and TrojDRL do drop off a decent amount from the original poisoning rate of $0.03\%$ presented in the main body, so it is possible they can find better performance as $\beta$ scales.
-Q-Incept was fairly consistent in ASR between $\beta_{low}$ and $\beta_{high}$ on all environments excluding CAGE-2.
-This large drop in ASR indicates that a poisoning rate of $0.5\%$ is too low for Q-Incept to be successful under bounded reward poisoning on CAGE-2.
-Across all environments and between $\beta_{low}$ and $\beta_{high}$ the BRR performance of each attack did not meaningfully change.
-A.3.2 Training Plots for Q*Bert and Safety Car
-Here in Figure 6 and Figure 8 we present the training curves for TrojDRL, Q-Incept, and SleeperNets on the Safety Car and Q*Bert environments respectively.
-We can see that all attacks perform similarly over time in terms of episodic return on both environments, but Q-Incept is the only attack to reach 100% ASR on average in both environments – doing so very quickly.
-Figure 6: Performance of TrojDRL, Q-Incept, and SleeperNets on the Safety Car environment in terms of ASR and episodic return.
-Figure 7: Performance of TrojDRL, Q-Incept, and SleeperNets on Q*Bert in terms of ASR and episodic return.
-A.3.3 How Often Does Q-Incept Alter the Agent’s Actions?
-Here we explore how often the Q-Incept attack alters chosen actions in the agent’s replay memory $\mathcal{D}$.
-In Figure 8 we see that the attack generally balances its action poisoning over time, altering actions on roughly 50% of the time steps it poisons.
-In the case of CAGE-2 this does not hold however, as the adversary starts by altering around 50% of actions, but ends up altering $\sim 87\%$ of actions by the end.
-To us this indicates that the difference in values between good and bad actions was much larger in CAGE-2 than in other environments, and furthermore that the agent was highly likely to choose these actions over others as training progressed.
-Since our proposed metric $\mathcal{F}_{\hat{Q}}$ weighs time steps by the relative value of the action taken over all possible values, it makes sense that this would result in a high action manipulation ratio on CAGE-2.
-Figure 8: Ratio of actions changed across poisoned states in all four environments.
-Values are measured as the number of actions changed divided by the total number of timesteps the attack has poisoned.
-A.4 Further Discussion
+A.3 Further Discussion
 In this section we provide further discussion on design choices made in this paper which were unable to fit in the main body.
-A.4.1 Motivation for Baselines
+A.3.1 Motivation for Baselines
 As mentioned in Section 6 we compare our Q-Incept attack against SleeperNets and TrojDRL as they represent the current state of the art for ubounded reward poisoning and forced action manipulation attacks respectively.
@@
 Trigger optimization is effective but orthogonal to the goals of this work as it can be generically applied to any attack.
-Furthermore, Rathbun et al.
-(2024) showed that BadRL without this trigger fine-tuning often performs worse than TrojDRL, likely since it uses methods to poison the most important – and thus hardest to poison – states in the MDP.
+Furthermore, (Rathbun et al., 2024) showed that BadRL without this trigger fine-tuning often performs worse than TrojDRL, likely since it uses methods to poison the most important – and thus hardest to poison – states in the MDP.
 Taking all of this into consideration we decided to omit BadRL from our empirical study.
 Therefore TrojDRL is the best baseline to use when comparing against static reward poisoning attacks using forced action manipulation.
-A.4.2 Motivation for Environments
-In this paper we study 4 environments – Q*Bert, Safety Car, CAGE-2, and Highway Merge.
-In the TrojDRL paper the authors focused their empirical studies towards Atari game tasks in the gym API Brockman et al.
-(2016).
+A.3.2 Motivation for Environments
+In this paper we study 4 environments – Q*Bert, Frogger, Safety Car, CAGE-2, and Highway Merge.
+In the TrojDRL paper the authors focused their empirical studies towards Atari game tasks in the gym API (Brockman et al., 2016).
 We think it is useful to include some of these environments like Q*Bert, as they are standard baselines for RL in general, however we believe it is critical to extend this study to further domains when studying the potential impacts of adversarial attacks.
-This belief is supported by the findings of Rathbun et al.
-(2024) who showed that TrojDRL – which consistently attains near 100% ASR on Atari environments without bounded reward poisoning constraints – often fails to achieve high ASR when tested on non-Atari environments.
+This belief is supported by the findings of (Rathbun et al., 2024) who showed that TrojDRL – which consistently attains near 100% ASR on Atari environments without bounded reward poisoning constraints – often fails to achieve high ASR when tested on non-Atari environments.
 Thus, to extend our study beyond the confines of Atari, we chose three other environments within the gymnasium API, allowing our code to work seamlessly between environments.
@@
 Lastly we selected Safety Car, also from the environments studied in SleeperNets, as it represents a simulation of real-world, robotic applications of RL and, similar to CAGE-2, uses non-image observations.
+A.3.3 Comments on Compute Overhead
+Training the Q-network used in Q-Incept does cause some computational overhead which may be detectable, but fortunately it isn’t too extreme.
+We ran tests on a desktop machine (2x RTX 4090, Threadripper 7980x) and found that SleeperNets, TrojDRL, and Q-Incept run at 1038, 987, and 730 simulation steps per second respectively against Atari Q*bert.
+However, our Q-network training runs in series with our PPO training, so it’s likely that significant performance increases can be found for Q-Incept by training in parallel.
+A.4 Further Experimental Results and Analysis
+A.4.1 Q-Incept Against DQN
+In Table 7 we present some preliminary results of Q-Incept against DQN.
+The results show that Q-Incept works well against both on policy (PPO) and off policy (DQN) methods.
+A better choice of metric $\mathcal{F}_{\hat{Q}}$ may be necessary to maximize performance against DQN, however.
+For this experiment we slightly alter the methodology of Q-Incept.
+Instead of poisoning data after each episode and storing it in the agent’s replay buffer, we instead poison data points after they’re sampled from the replay buffer.
+The reason for this is that the agent’s policy along with the Q-Network of Q-Incept change throughout training, therefore datapoints that were poisoned early in training may become less effective as training progresses.
+In the case of PPO this wasn’t an issue, since trajectories are only used once and then discarded.
+For DQN, however, trajectories are stored over a longer time horizon during training, so we need to account for this as an attacker.
+Equivalently the attacker could keep track of all the data points it has poisoned and properly update them over time.
+Beta ASR StDev(ASR) BR StDev(BR)
+0% (No Poisoning) N/a N/a 13,724 974
+0.5% 92.6% 4.1% 14,092 1,127
+1.0% 88.5% 3.4% 13,574 910
+Table 7: Performance of Q-Incept against DQN agents training on Atari Q*bert. Results are averaged over 5 runs.
+A.4.2 Q-Incept at Different Poisoning Rates
+In this section we evaluate the Attack Success Rate performance of Q-Incept at lower $\beta$ values along with the stability of the agent’s Benign Return at higher $\beta$ values.
+First, in Table 8 we evaluate Q-Incept on Atari Q*bert at $\beta$ values lower than those presented in the main body of the paper.
+We can see that even at a poisoning rate as low as $\beta=0.03\%$ Q-Incept is still able to achieve near 100% ASR.
+This is in line with the performance of methods like SleeperNets which presented similar results at $\beta=0.03\%$ while operating under unconstrained rewards.
+$\beta=0.03\%$ seems to be the minimum necessary poisoning rate for Q-Incept on Q*bert however, as the attack success rate drops off dramatically at $\beta=0.01\%$.
+Beta ASR StDev(ASR) BR StDev(BR)
+0.3% 100% 0% 18,381 882
+0.1% 100% 0% 17,749 1,380
+0.05% 100% 0% 17,937 1,304
+0.03% 98.3% 2.9% 16,573 873
+0.01% 21.1% 6.2% 16,374 2,088
+Table 8: Performance of Q-Incept at lower poisoning rates against a PPO agent training on Q*bert.
+In Table 9 we compare Q-Incept to the baselines on the CAGE-2 environment to study their stability at higher $\beta$ values.
+We see that Q-Incept is the only method that improves in ASR as $\beta$ increases, and is also the most stable in terms of BR – never falling below -50.98%.
+In contrast, SleeperNets and TrojDRL are both highly inconsistent in terms of BR – falling to $-57.82$ and $-64.41$, respectively – while also failing to achieve an ASR above $6\%$, even at $\beta=2\%$.
+$\beta$ 0.5% 1% 1.5% 2%
+Metric ASR $\sigma$ ASR $\sigma$ ASR $\sigma$ ASR $\sigma$
+Q-Incept 30.06% 15.01% 93.21% 15.13% 100% 0.00% 98.62% 2.14%
+SleeperNets 0.00% 0.00% 0.06% 0.12% 0.54% 0.89% 1.86% 3.21%
+TrojDRL 0.00% 0.00% 5.64% 7.73% 2.24% 3.74% 5.11% 4.99%
+Metric BR $\sigma$ Br $\sigma$ BR $\sigma$ BR $\sigma$
+Q-Incept -48.64 16.68 -45.49 8.04 -50.98 7.37 -49.41 7.50
+SleeperNets -57.82 8.38 -50.18 8.52 -41.44 11.59 -50.67 9.16
+TrojDRL -39.28 9.60 -53.11 8.71 -52.77 13.54 -64.41 10.63
+Table 9: Comparison of Q-Incept, SleeperNets, and TrojDRL on CAGE at different values of $\beta$.
+A.4.3 Q-Incept with an Oracle Q-Network
+We noticed that Highway Merge was the only environment on which Q-Incept was unable to attain an average ASR above 90%, leading us to question if our Q-function based approach was incorrect or if our online DQN approximation $\hat{Q}$ wasn’t converging quickly enough.
+To test this we devised Oracle-Incept – which uses an oracle Q-function pre-trained with DQN until convergence – as a hypothetical, stronger attack by an adversary with direct access to the benign MDP.
+In Figure 7 we can see that Oracle-Incept improves greatly over Q-Incept, reaching an average ASR of 93.38%.
+This indicates that better Q-function approximations lead to better performance - validating that both our chosen metric and attack approach scale properly with the accuracy of $\hat{Q}$.
+Thus, adversaries capable of using Q-function estimations with faster convergence can expect greater attack success.
+Figure 7: Comparison between Q-Incept, Oracle-Incept, TrojDRL, and SleeperNets on Highway Merge at $\beta=10\%$. We can see that the Oracle-Incept shows significant improvements in ASR.
+A.4.4 Training Plots for Remaining Environments
+Here in Figure 8, Figure 9, Figure 10 and Figure 11 we present the training curves for TrojDRL, Q-Incept, and SleeperNets on the Frogger, Pacman, Breakout and Safety Car environments respectively.
+Plots for Q*Bert are found in Figure 4, CAGE-2 can be found in Figure 3, and Highway Merge in Figure 7.
+We can see that all attacks perform similarly over time in terms of episodic return on both environments, but Q-Incept is the only attack to reach 100% ASR on average in both environments – doing so very quickly.
+Figure 8: Performance of TrojDRL, Q-Incept, and SleeperNets on Frogger at $\beta=0.3\%$
+Figure 9: Performance of TrojDRL, Q-Incept, and SleeperNets on Pacman at $\beta=0.3\%$
+Figure 10: Performance of TrojDRL, Q-Incept, and SleeperNets on Breakout at $\beta=0.3\%$
+Figure 11: Performance of TrojDRL, Q-Incept, and SleeperNets on the Safety Car environment at $\beta=0.1\%$
+A.4.5 How Often Does Q-Incept Alter the Agent’s Actions?
+Here we explore how often Q-Incept chosen actions in the agent’s replay memory $\mathcal{D}$.
+We measure this ratio as the number of actions changed divided by the total number of timesteps the attack has poisoned.
+In Figure 12 and Figure 13 we see that the attack generally balances its action poisoning over time, altering actions on roughly 50% of the time steps it poisons.
+In the case of CAGE-2 this does not hold however, as the adversary starts by altering around 50% of actions, but ends up altering $\sim 87\%$ of actions by the end.
+To us this indicates that the difference in values between good and bad actions was much larger in CAGE-2 than in other environments, and furthermore that the agent was highly likely to choose these actions over others as training progressed.
+Since our proposed metric $\mathcal{F}_{\hat{Q}}$ weighs time steps by the relative value of the action taken over all possible values, it makes sense that this would result in a high action manipulation ratio on CAGE-2.
+Figure 12: Ratio of actions changed across poisoned states in our four Atari environments – (from top left to bottom right) Q*bert, Frogger, Pacman, and Breakout.
+Values are measured as the ratio of actions changed to the total number of poisoned timesteps.
+Figure 13: Ratio of actions changed across poisoned states in our three non-Atari environments – (frop top left to bottom) CAGE, Highway Merge, and Safety Car.
+Values are measured as the ratio of actions changed to the total number of poisoned timesteps.